Theory and Applications of Categories, Vol. 18, No. 15, 2007, pp. 400–439.

ENRICHED MODEL CATEGORIES AND AN APPLICATION TO ADDITIVE ENDOMORPHISM SPECTRA

DANIEL DUGGER AND BROOKE SHIPLEY

Abstract. We define the notion of an additive model category and prove that any stable, additive, combinatorial model category ℳ has a model enrichment over Sp^{Σ}(s𝒜b) (symmetric spectra based on simplicial abelian groups). So to any object X ∈ ℳ one can attach an endomorphism ring object, denoted hEnd_{ad}(X), in the category Sp^{Σ}(s𝒜b). We establish some useful properties of these endomorphism rings. We also develop a new notion in enriched category theory which we call ‘adjoint modules’. This is used to compare enrichments over one symmetric monoidal model category with enrichments over a Quillen equivalent one. In particular, it is used here to compare enrichments over Sp^{Σ}(s𝒜b) and chain complexes.

1. Introduction

A model category is called additive if two conditions are satisfied. First, its hom-sets must have natural structures of abelian groups with respect to which composition is biadditive. Secondly, the abelian group structures on these hom-sets must interact well with the notion of ‘higher homotopies’. We give a precise definition in Section 6. Examples of additive model categories include chain complexes over a ring and differential graded modules over a differential graded algebra, as one should expect.

Recall that a category is locally presentable if it is cocomplete and all objects are small with respect to large enough filtered colimits; see [AR]. A model category is called combinatorial if it is cofibrantly-generated and its underlying category is locally presentable. Finally, a model category is stable if it is pointed and the suspension functor is an auto-equivalence of the homotopy category.

In [D4] it was shown that any stable, combinatorial model category can be naturally enriched over the category Sp^{Σ} of symmetric spectra. This enrichment is invariant under Quillen equivalences in a certain sense. In the present paper we extend the results of [D4] to the additive case. We show that any stable, combinatorial, additive model category has a natural enrichment over Sp^{Σ}(s𝒜b)—the category of symmetric spectra based on simplicial abelian groups. This enrichment is not an invariant of Quillen equivalence, however: if ℳ and 𝒩 are stable, combinatorial, additive model categories which are connected by a zig-zag of Quillen equivalences (with the intermediate model categories not necessarily additive) then the additive enrichments for ℳ and 𝒩 need not be related. We can prove that the enrichments are equivalent, however, if all the intermediate steps in the zig-zag are additive.

The first author was partially supported by NSF grant DMS0604354. The second author was partially supported by NSF grants DMS0417206 and DMS0706877.

Received by the editors 2007-02-15 and, in revised form, 2007-07-15.

Transmitted by Ieke Moerdijk. Published on 2007-09-09.

2000 Mathematics Subject Classification: 18D20, 55U35, 55P42, 18E05.

Key words and phrases: model categories, symmetric spectra, endomorphism ring.

\begin{matrix}c \\ ◯ \end{matrix} Daniel Dugger and Brooke Shipley, 2007. Permission to copy for private use granted.

1.1. Remark. The tools developed in this paper are applied in [DS2]. Two additive model categories ℳ and 𝒩 are called additively Quillen equivalent if there is a zigzag of Quillen equivalences between ℳ and 𝒩 in which every intermediate model category is additive. It is a strange fact, established in [DS2], that additive model categories can be Quillen equivalent but not additively Quillen equivalent. The demonstration of this fact uses the model enrichments developed in the present paper.

We should explain up front that there are really three separate things going on in this paper. One is the development of the theory of additive model categories, taken up in Sections 6 and 7. The second is the construction of the model enrichment over Sp^{Σ}(s𝒜b), which is begun in Section 8. Most of the details of the model enrichment exactly follow the pattern in [D4]. There is one extra result we wish to consider, though, which involves comparing model enrichments over Sp^{Σ}(s𝒜b) to model enrichments over the Quillen equivalent category Ch of chain complexes of abelian groups. For this last issue we need to develop quite a bit more about enriched model categories than is available in the literature. Since this foundational material is important in its own right, we include it at the very beginning as Sections 2 through 5.

1.2. A closer look at the results. To describe the results in more detail we need to recall some enriched model category theory; specifically, we need the notions of model enrichment and quasi-equivalence from [D4]. Let ℳ be a model category and 𝒱 be a symmetric monoidal model category. Briefly, a model enrichment is a bifunctor τ: ℳ^{op} × ℳ → 𝒱 together with composition maps τ(Y, Z) ⊗ τ(X, Y) → τ(X, Z) which are associative and unital. The bifunctor must interact well with the model category structure—see [D4] for an explicit list of the necessary axioms, or Section 2.3 for a summary.

There is a notion of when two model enrichments of ℳ by 𝒱 are ‘quasi-equivalent’, which implies that they carry the same homotopical information. This takes longer to describe, but the reader can again find it in Section 2.3. We let ME_{0}(ℳ, 𝒱) denote the quasi-equivalence classes of model enrichments.

If L: ℳ ⇄ 𝒩: R is a Quillen pair, there are induced functors L_{∗}: ME_{0}(ℳ, 𝒱) → ME_{0}(𝒩, 𝒱) and L^{∗}: ME_{0}(𝒩, 𝒱) → ME_{0}(ℳ, 𝒱). When (L, R) is a Quillen equivalence these are inverse bijections.

Using the above language, we can state the basic results. These are proved in Sections 8 and 9.

1.3. Theorem. If ℳ is a stable, additive, combinatorial model category, then there is a canonical element σ_{ℳ} ∈ ME_{0}(ℳ, Sp^{Σ}(s𝒜b)). If L: ℳ → 𝒩 is a Quillen equivalence then L_{∗}(σ_{ℳ} ) = σ_{𝒩} and L^{∗}(σ_{𝒩} ) = σ_{ℳ} .

If X ∈ ℳ, choose a cofibrant-fibrant object X̂ which is weakly equivalent to X. Also, choose a specific model enrichment σ_{ℳ} representing the equivalence class in the above theorem. Then σ_{ℳ} (X̂, X̂) gives a ring object in Sp^{Σ}(s𝒜b). The resulting isomorphism class in the homotopy category Ho (ℛing[Sp^{Σ}(s𝒜b)]) only depends on the homotopy type of X, not on the choice of X̂ or the representative σ_{ℳ} . We write hEnd_{ad}(X) for any ring object in this isomorphism class, and call it the additive homotopy endomorphism spectrum of X.

1.4. Proposition. Let ℳ and 𝒩 be additive, stable, combinatorial model categories. Suppose ℳ and 𝒩 are Quillen equivalent through a zig-zag of additive (but not necessarily combinatorial) model categories. Let X ∈ ℳ, and let Y ∈ 𝒩 correspond to X under the derived equivalence of homotopy categories. Then hEnd_{ad}(X) and hEnd_{ad}(Y) are weakly equivalent in ℛing[Sp^{Σ}(s𝒜b)].

Any ring object R in Sp^{Σ}(s𝒜b) gives rise to a ring object in Sp^{Σ} by forgetting the abelian group structure—this is called the Eilenberg-Mac Lane spectrum associated to R. Recall that in [D4] it was shown how to attach to any X in a stable, combinatorial model category an isomorphism class in Ho (ℛing[Sp^{Σ}]). This was called the homotopy endomorphism spectrum of X, and denoted hEnd(X). We have the following:

1.5. Proposition. Given X ∈ ℳ as above, the homotopy endomorphism spectrum hEnd(X) is the Eilenberg-Mac Lane spectrum associated to hEnd_{ad}(X).

Finally, we have two results explaining how to compute hEnd_{ad}(X) when the model category ℳ has some extra structure. Recall that if 𝒞 is a symmetric monoidal model category then a 𝒞-model category is a model category equipped with compatible tensors, cotensors, and enriched hom-objects over 𝒞 satisfying the analogue of SM7. See Section 2 for more detailed information. For X, Y ∈ ℳ we denote the enriched hom-object by \underline{ℳ}_{𝒞}(X, Y).

Note that a Sp^{Σ}(s𝒜b)-model category is automatically additive and stable. This follows from Corollary 6.9 below, and the appropriate analogue of [SS2, 3.5.2] or [GS, 3.2].

1.6. Proposition. Let ℳ be a combinatorial Sp^{Σ}(s𝒜b)-model category. Let X ∈ ℳ be cofibrant-fibrant. Then hEnd_{ad}(X) is weakly equivalent to the enriched hom-object \underline{ℳ}_{Sp^{Σ}(s𝒜b)} (X, X).

In [S] it is shown that the model categories of rings in Sp^{Σ}(s𝒜b) and in Ch are Quillen equivalent. This is recalled in Section 9. Note that the rings in the category Ch are just differential graded algebras (dgas). The associated derived functors will be denoted H′: 𝒟G𝒜 ⇄ ℛingSp^{Σ}(s𝒜b): Θ′. (The reason for the ‘primes’ is that in [S] the functors H and Θ are functors between 𝒟G𝒜 and Hℤ-algebras with ℛingSp^{Σ}(s𝒜b) an intermediate category.) We then define the homotopy endomorphism dga of X to be Θ′[hEnd_{ad}(X)] and write hEnd_{dga}(X). Obviously, this carries exactly the same information as hEnd_{ad}(X). In fact, H′[hEnd_{dga}(X)] is weakly equivalent to hEnd_{ad}(X) since H′ and Θ′ are inverse equivalences on the homotopy category level.

As above, we remark that a Ch-model category is automatically additive and stable, by Corollary 6.9 and the appropriate analogue of [SS2, 3.5.2].

1.7. Proposition. Let ℳ be a combinatorial Ch-model category. Assume ℳ has a generating set of compact objects, as defined in (5.1) below. Let X ∈ ℳ be cofibrant-fibrant. Then \underline{ℳ}_{Ch}(X, X) is weakly equivalent to hEnd_{dga}(X).

The assumption about the generating set in the above proposition is probably unnecessary, but we don’t know how to remove it. It is satisfied in most cases of interest.

1.8. Remark. The study of dg-categories seems to be of current interest—see, for example, [Dr, T]. A dg-category is simply a category enriched over unbounded chain complexes Ch_{k}, where k is some commutative ground ring. We remark that the homotopy theory of dg-categories over ℤ is essentially the same as that of Sp^{Σ}(s𝒜b)-categories (this follows from results of [S] and [SS3]). So the present paper may be regarded as attaching to any stable, combinatorial, additive model category an associated dg-category.

The proof of Proposition 1.7 is not hard, but it requires a careful comparison of enrichments over Ch and Sp^{Σ}(s𝒜b). This reduces to an abstract problem in enriched model category theory, but the necessary tools do not seem to be available in the literature. The first part of the paper is spent developing them.

We can briefly describe the main issue that arises. Suppose that 𝒞 and 𝒟 are symmetric monoidal categories, and that there is an adjoint pair F: 𝒞 ⇄ 𝒟: G in which the right adjoint G is lax monoidal. Also suppose that ℳ is a 𝒞-module category, 𝒩 is a 𝒟-module category, and that

L: ℳ ⇄ 𝒩: R

is another adjoint pair (all the terminology is defined later in the paper).

Then for any object X ∈ 𝒩 we may consider the endomorphism monoid \underline{𝒩}_{𝒟} (X, X). This is a monoid in 𝒟, and applying G to it gives a monoid in 𝒞. One can also consider the endomorphism monoid \underline{ℳ}_{𝒞}(RX, RX), which is another monoid in 𝒞. In order to compare these two monoids, one needs some compatibility conditions between the adjoint pairs (F, G) and (L, R). This leads to our definition of what we call an adjoint module. The basic theory of such things is developed in Sections 3–4. This notion has other applications, most notably in [GS].

1.9. Organization of the paper. Section 2 recalls the basics of enriched model category theory as used in [D4]. The new work begins in Sections 3 and 4 where we develop the notion of adjoint modules. This is used in Section 5 to prove a technical theorem about transporting enrichments over one symmetric monoidal model category to a Quillen equivalent one. Sections 6 through 9 contain the main results on additive model categories and Sp^{Σ}(s𝒜b)-enrichments. Appendix A reviews and expands on material from [SS3], which is needed in Section 4.9.

1.10. Notation and terminology. This paper is a sequel to [D4], and we will assume the reader is familiar with the machinery developed therein. In particular, we assume a familiarity with model enrichments and quasi-equivalences; see Section 2.3 for quick summaries, though. If ℳ and 𝒩 are model categories then by a Quillen map L: ℳ → 𝒩 we mean an adjoint pair of Quillen functors L: ℳ ⇄ 𝒩: R, where L is the left adjoint. Also, if 𝒞 is a category then we’ll write 𝒞(X, Y) for the set of maps from X to Y in 𝒞. If 𝒞 is a symmetric monoidal category, then the monoids in 𝒞 are sometimes called “monoids” and sometimes called “rings”—we use these terms interchangeably.

2. Enriched model categories

In this section we review the notion of a model category ℳ being enriched over a second model category 𝒞. This situation comes in two varieties. If for every two objects X, Y ∈ ℳ one has a ‘mapping object’ \underline{ℳ}_{𝒞}(X, Y) in 𝒞 together with composition maps (subject to certain axioms), then this is called a model enrichment. If for every X ∈ ℳ and c ∈ 𝒞 one also has objects X ⊗ c and F (c, X) in ℳ, related by adjunctions to the mapping objects and also subject to certain axioms, then we say that ℳ is a 𝒞-model category. Thus, a 𝒞-model category involves a model enrichment plus extra data.

There are two main examples to keep in mind. A simplicial model category is just another name for an s𝒮et-model category. And if ℳ is any model category, then the hammock localization of Dwyer-Kan [DK] is an example of a model enrichment of ℳ over s𝒮et.

2.1. Symmetric monoidal model categories. Let 𝒞 be a closed symmetric monoidal category. This says that we are given a bifunctor ⊗, a unit object 1_{𝒞}, together with associativity, commutativity, and unital isomorphisms making certain diagrams commute (see [Ho1, Defs. 4.1.1, 4.1.4] for a nice summary). The ‘closed’ condition says that there is also a bifunctor (a, b) ↦ \underline{𝒞}(a, b) ∈ 𝒞 together with a natural isomorphism

𝒞(a, \underline{𝒞}(b, c)) ≅ 𝒞(a ⊗ b, c).

Note that this gives isomorphisms 𝒞(1_{𝒞}, \underline{𝒞}(a, b)) ≅ 𝒞(1_{𝒞} ⊗ a, b) ≅ 𝒞(a, b).

A symmetric monoidal model category consists of a closed symmetric monoidal category 𝒞, together with a model structure on 𝒞, satisfying two conditions:

(1) The analogue of SM7, as given in either [Ho1, 4.2.1] or [Ho1, 4.2.2(2)].

(2) A unit condition given in [Ho1, 4.2.6(2)].

2.2. 𝒞-model categories. Let 𝒞 be a symmetric monoidal category. One defines a closed 𝒞-module category to be a category ℳ equipped with natural constructions which assign to every X, Z ∈ ℳ and c ∈ 𝒞 objects

X ⊗ c ∈ ℳ,  F (c, Z) ∈ ℳ,  and \underline{ℳ}_{𝒞}(X, Z) ∈ 𝒞.

One requires, first, that there are natural isomorphisms (X ⊗ a) ⊗ b ≅ X ⊗ (a ⊗ b) and X ⊗ 1_{𝒞} ≅ X making certain diagrams commute (see [Ho1, Def. 4.1.6]). One of these diagrams is a pentagon for four-fold associativity. We also require natural isomorphisms

ℳ(X ⊗ a, Z) ≅ ℳ(X, F (a, Z)) ≅ 𝒞(a, \underline{ℳ}_{𝒞}(X, Z))  (2.2)

(see [Ho1, 4.1.12]).

Finally, suppose 𝒞 is a symmetric monoidal model category. A 𝒞-model category is a model category ℳ which is also a closed 𝒞-module category and where the two conditions from [Ho1, 4.2.18] hold: these are again the analogue of SM7 and a unit condition.

2.3. Model enrichments. Let ℳ be a model category and let 𝒞 be a symmetric monoidal model category. Recall from [D4, 3.1] that a model enrichment of ℳ by 𝒞 is a bifunctor σ: ℳ^{op} × ℳ → 𝒞 which is equipped with composition pairings σ(Y, Z) ⊗ σ(X, Y) → σ(X, Z) and unit maps 1_{𝒞} → σ(X, X) satisfying associativity and unital conditions. There is also a compatibility condition between the functor structure and the unit maps. Finally, one assumes that if X → X′ is a weak equivalence between cofibrant objects and Y → Y′ is a weak equivalence between fibrant objects then the maps σ(X, Y) → σ(X, Y′) and σ(X′, Y) → σ(X, Y) are weak equivalences. See [D4, Section 3.1].

There is a notion of quasi-equivalence encoding when two model enrichments are ‘the same’. This is also given in [D4, Section 3.1]. To define this we need two preliminary notions.

Let σ and τ be two model enrichments of ℳ by 𝒞. By a σ − τ bimodule we mean a collection of objects M(a, b) ∈ 𝒞 for every a, b ∈ 𝒞, together with multiplication maps

σ(b, c) ⊗ M(a, b) → M(a, c) and M(b, c) ⊗ τ(a, b) → M(a, c)

which are natural in a and c. Associativity and unital conditions are again assumed, although we will not write these down. One also requires that for any a, b, c, d ∈ 𝒞 the two obvious maps

σ(c, d) ⊗ M(b, c) ⊗ τ(a, b) ⇉ M(a, d)

are equal.

It is perhaps not quite obvious, but M becomes a bifunctor via the multiplication maps from σ and τ and the fact that σ and τ are bifunctors. See [D4, Section 2.2].

A pointed σ − τ bimodule is a bimodule M together with a collection of maps 1_{𝒞} → M(c, c) for every c ∈ 𝒞, such that for any map a → b the square

\xymatrix{1_{𝒞} \ar[r] \ar[d] & M(a, a) \ar[d] \\ M(b, b) \ar[r] & M(a, b)}

commutes.

A quasi-equivalence between two model enrichments σ and τ consists of a pointed σ − τ bimodule M such that the compositions

σ(a, b) ⊗ 1_{𝒞} → σ(a, b) ⊗ M(a, a) → M(a, b) and
1_{𝒞} ⊗ τ(a, b) → M(b, b) ⊗ τ(a, b) → M(a, b)

are weak equivalences whenever a is cofibrant and b is fibrant.

The notion of quasi-equivalence generates an equivalence relation on the class of model enrichments of ℳ by 𝒞. We write ME_{0}(ℳ, 𝒞) for the collection of equivalence classes of model enrichments. When we say that two enrichments σ and τ are ‘quasi-equivalent’ we mean that they are in the same equivalence class; note that this means there is a chain of model enrichments σ = σ_{1}, σ_{2}, …, σ_{n} = τ and pointed σ_{i} − σ_{i+1} bimodules M_{i} giving quasi-equivalences between each step in the chain.

If L: ℳ → 𝒩 is a Quillen map then by [D4, Prop. 3.14] there are induced maps L_{∗}: ME_{0}(ℳ, 𝒞) → ME_{0}(𝒩, 𝒞) and L^{∗}: ME_{0}(𝒩, 𝒞) → ME_{0}(ℳ, 𝒞). When L is a Quillen equivalence these are inverse bijections.

2.4. Monoidal functors. Suppose that 𝒞 and 𝒟 are symmetric monoidal model categories, and that F: 𝒞 ⇄ 𝒟: G is a Quillen pair.

First of all, recall that G is called lax monoidal if there is a natural transformation

G(X) ⊗ G(Y) → G(X ⊗ Y)

and a map 1_{𝒞} → G(1_{𝒟} ) which are compatible with the associativity and unital isomorphisms in 𝒞 and 𝒟. A lax monoidal functor takes monoids in 𝒟 to monoids in 𝒞.

A lax monoidal functor is called strong monoidal if the above maps are actually isomorphisms.

If G is lax monoidal then the adjunction gives rise to induced maps F (1_{𝒞}) → 1_{𝒟} and F (A ⊗ B) → F (A) ⊗ F (B). Following [SS3, Section 3], we say that (F, G) is a weak monoidal Quillen equivalence if G is lax monoidal and two extra conditions hold. First, for some cofibrant replacement A → 1_{𝒞}, the induced map F (A) → F (1_{𝒞}) → 1_{𝒟} is a weak equivalence. Second, for any two cofibrant objects A, B ∈ 𝒞 the map F (A ⊗ B) → F (A) ⊗ F (B) is a weak equivalence.

3. Adjoint modules

In this section and the next we deal with the general situation of one Quillen pair enriched over another Quillen pair. Let 𝒞 and 𝒟 be symmetric monoidal model categories, let ℳ be a 𝒞-model category, and let 𝒩 be a 𝒟-model category. Let

F: 𝒞 ⇄ 𝒟: G and L: ℳ ⇄ 𝒩: R

be two Quillen pairs, where we assume that G is lax monoidal (see Section 2.4). As usual, we’ll write \underline{ℳ}_{𝒞}(X, Y) and \underline{𝒩}_{𝒟} (X, Y) for the enriched morphism objects over 𝒞 and 𝒟, respectively.

Finally, let Y be a cofibrant-fibrant object in 𝒩. Then \underline{𝒩}_{𝒟} (Y, Y) is a monoid in 𝒟, and so G(\underline{𝒩}_{𝒟} (Y, Y)) is a monoid in 𝒞. Alternatively, we may choose a cofibrant-replacement 𝒬RY −↠^{∼} RY and consider the 𝒞-monoid \underline{ℳ}_{𝒞}(𝒬RY, 𝒬RY). How can we compare these two monoids, and under what conditions will they be weakly equivalent?

This question can be answered by requiring certain compatibility conditions between (L, R) and (F, G). The goal of the present section is to write down these conditions; this culminates in Definition 3.8, where we define what it means for (L, R) to be an adjoint module over (F, G). The next section uses this to tackle the problem of comparing enrichments.

3.1. Compatibility structure. Before we can develop the definition of an adjoint module we need the following statement. For the moment we only assume that (F, G) and (L, R) are adjunctions. That is, we temporarily drop the assumptions that they are Quillen pairs and that G is lax monoidal.

3.2. Proposition. There is a canonical bijection between natural transformations of the following four types:

(i) G\underline{𝒩}_{𝒟} (LX, Y) → \underline{ℳ}_{𝒞}(X, RY)

(ii) L(X ⊗ c) → LX ⊗ F c

(iii) RY ⊗ Gd → R(Y ⊗ d)

(iv) G\underline{𝒩}_{𝒟} (X, Y) → \underline{ℳ}_{𝒞}(RX, RY).

Proof. This is a routine exercise in adjunctions. We will only do some pieces of the argument and leave the rest to the reader.

Suppose given a natural transformation G\underline{𝒩}_{𝒟} (LX, Y) → \underline{ℳ}_{𝒞}(X, RY). For any c ∈ 𝒞 one therefore has the composition

\xymatrix{ & 𝒞(c, G\underline{𝒩}_{𝒟} (LX, Y)) \ar[d]^{≅} \ar[r] & 𝒞(c, \underline{ℳ}_{𝒞}(X, RY)) \ar[d]^{≅} \\ 𝒩(LX ⊗ F c, Y) \ar[r]^{≅} & 𝒟(F c, \underline{𝒩}_{𝒟} (LX, Y)) & ℳ(X ⊗ c, RY) \ar[r]^{≅} & 𝒩(L(X ⊗ c), Y).} (3.3)

By the Yoneda Lemma this gives a map L(X ⊗ c) → LX ⊗ F c, and this is natural in both X and c.

Likewise, suppose given a natural transformation L(X ⊗ c) → LX ⊗ F c. Then for Y ∈ 𝒩 and d ∈ 𝒟 we obtain

L(RY ⊗ Gd) → LRY ⊗ F Gd → Y ⊗ d

where the second map uses the units of the adjunctions. Taking the adjoint of the composition gives RY ⊗ Gd → R(Y ⊗ d), as desired.

Finally, suppose again that we have a natural transformation G\underline{𝒩}_{𝒟} (LX, Y) → \underline{ℳ}_{𝒞}(X, RY). For X, Y ∈ 𝒩 consider the composite

G\underline{𝒩}_{𝒟} (X, Y) → G\underline{𝒩}_{𝒟} (LRX, Y) → \underline{ℳ}_{𝒞}(RX, RY)

where the first map is obtained by applying G to \underline{𝒩}_{𝒟} (X, Y) → \underline{𝒩}_{𝒟} (LRX, Y) induced by the unit LRX → X. The above composite is our natural transformation of type (iv).

We have constructed maps (i) → (ii), (ii) → (iii), and (i) → (iv). We leave it to the reader to construct maps in the other directions and verify that one obtains inverse bijections.

3.4. Remark. Suppose we are given a natural transformation γ: G\underline{𝒩}_{𝒟} (LX, Y) → \underline{ℳ}_{𝒞}(X, RY). Using the bijections from the above result, we obtain natural transformations of types (ii), (iii), and (iv). We will also call each of these γ, by abuse.

The next proposition lists the key homotopical properties required for (L, R) to be a Quillen adjoint module over (F, G).

3.5. Proposition. Assume that (F, G) and (L, R) are Quillen pairs and that we have a natural transformation γ: G\underline{𝒩}_{𝒟} (LX, Y) → \underline{ℳ}_{𝒞}(X, RY).

(a) The following two conditions are equivalent:

• The map γ: G\underline{𝒩}_{𝒟} (LX, Y) → \underline{ℳ}_{𝒞}(X, RY) is a weak equivalence whenever X is cofibrant and Y is fibrant.

• The map γ: L(X ⊗ c) → LX ⊗ F c is a weak equivalence whenever X and c are both cofibrant.

(b) If (L, R) is a Quillen equivalence, the conditions in (a) are also equivalent to:

• For any cofibrant replacement 𝒬RX → RX, the composite map

G\underline{𝒩}_{𝒟} (X, Y) ⟶^{γ} \underline{ℳ}_{𝒞}(RX, RY) → \underline{ℳ}_{𝒞}(𝒬RX, RY)

is a weak equivalence whenever X is cofibrant-fibrant and Y is fibrant.

(c) Assume that both (L, R) and (F, G) are Quillen equivalences. Then the conditions in (a) and (b) are also equivalent to:

• For any cofibrant replacements 𝒬RY → RY and 𝒬′Gd → Gd and any fibrant replacement Y ⊗ d → ℱ(Y ⊗ d), the composite

𝒬RY ⊗ 𝒬′Gd → RY ⊗ Gd ⟶^{γ} R(Y ⊗ d) → Rℱ(Y ⊗ d)

is a weak equivalence whenever Y and d are cofibrant and fibrant.

Proof. This is routine and basically follows from the adjunctions in Proposition 3.2 with the following two additions. For the equivalence in part (a), consider the maps from 3.3 in the respective homotopy categories. For the equivalence with (b), note that the composite in (b) agrees with the composite

G\underline{𝒩}_{𝒟} (X, Y) → G\underline{𝒩}_{𝒟} (L𝒬RX, Y) ⟶^{γ} \underline{ℳ}_{𝒞}(𝒬RX, RY).

The above homotopical properties need to be supplemented by categorical associativity and unital properties which are listed in the next two propositions. Then, after stating these categorical properties, we finally state the definition of a Quillen adjoint module.

3.6. Proposition. Assume G is lax monoidal. Note that this gives a lax comonoidal structure on F, by adjointness. Let γ again denote a set of four corresponding natural transformations of types (i)–(iv). Then the conditions in (a) and (b) below are equivalent:

(a) The diagrams

\xymatrix{L((X ⊗ c) ⊗ c′) \ar[r]^{γ} \ar[d]^{≅} & L(X ⊗ c) ⊗ F c′ \ar[r]^{γ⊗1} & (LX ⊗ F c) ⊗ F c′ \ar[d]^{≅} \\ L(X ⊗ (c ⊗ c′)) \ar[r]^{γ} & LX ⊗ F (c ⊗ c′) \ar[r] & LX ⊗ (F c ⊗ F c′)}

all commute, for any X, c, c′.

(b) The diagrams

\xymatrix{RY ⊗ (Gd ⊗ Gd′) \ar[r] \ar[d]^{≅} & RY ⊗ G(d ⊗ d′) \ar[r]^{γ} & R(Y ⊗ (d ⊗ d′)) \ar[d]^{≅} \\ (RY ⊗ Gd) ⊗ Gd′ \ar[r]^{γ⊗1} & R(Y ⊗ d) ⊗ Gd′ \ar[r]^{γ} & R((Y ⊗ d) ⊗ d′)}

all commute, for any Y, d, d′.

If G is lax symmetric monoidal, then the above (equivalent) conditions imply the following one:

(c) The diagrams

[figure: G\underline{𝒩}_{𝒟} (Y, Z) ⊗ G\underline{𝒩}_{𝒟} (X, Y); γ⊗γ; G(\underline{𝒩}_{𝒟} (Y, Z) ⊗ \underline{𝒩}_{𝒟} (X, Y)); G\underline{𝒩}_{𝒟} (X, Z); γ; \underline{ℳ}_{𝒞}(RY, RZ) ⊗ \underline{ℳ}_{𝒞}(RX, RY); \underline{ℳ}_{𝒞}(RX, RZ)]

commute for any X, Y, and Z.

Proof. The equivalence of (a) and (b) is extremely tedious but routine; we leave it to the reader. For (c), note that by using the adjunction 𝒞(c, \underline{ℳ}_{𝒞}(RX, RZ)) ≅ ℳ_{𝒞}(RX ⊗ c, RZ) the two ways of going around the diagram correspond to two maps

RX ⊗ [G\underline{𝒩}_{𝒟} (Y, Z) ⊗ G\underline{𝒩}_{𝒟} (X, Y)] ⟶ RZ.

One of these is the composite

\xymatrix{ & RX ⊗ [G\underline{𝒩}_{𝒟} (Y, Z) ⊗ G\underline{𝒩}_{𝒟} (X, Y)] \ar[r] & RX ⊗ G[\underline{𝒩}_{𝒟} (Y, Z) ⊗ \underline{𝒩}_{𝒟} (X, Y)] \ar[d] \\ RZ & R(X ⊗ \underline{𝒩}_{𝒟} (X, Z)) \ar[l] & RX ⊗ G[\underline{𝒩}_{𝒟} (X, Z)] \ar[l]^{γ}}

The other is the composite

\xymatrix{RX ⊗ [G\underline{𝒩}_{𝒟} (Y, Z) ⊗ G\underline{𝒩}_{𝒟} (X, Y)] \ar[r]^{≅} & RX ⊗ [G\underline{𝒩}_{𝒟} (X, Y) ⊗ G\underline{𝒩}_{𝒟} (Y, Z)] \ar[d] \\ [R(X ⊗ \underline{𝒩}_{𝒟} (X, Y))] ⊗ G\underline{𝒩}_{𝒟} (Y, Z) \ar[d] & [RX ⊗ G\underline{𝒩}_{𝒟} (X, Y)] ⊗ G\underline{𝒩}_{𝒟} (Y, Z) \ar[l]_{γ⊗1} \\ RY ⊗ G\underline{𝒩}_{𝒟} (Y, Z) \ar[r]^{γ} & R(Y ⊗ \underline{𝒩}_{𝒟} (Y, Z)) \ar[r] & RZ.}

The commutativity isomorphism comes into the first stage of this composite because of how the composition map \underline{ℳ}_{𝒞}(RY, RZ) ⊗ \underline{ℳ}_{𝒞}(RX, RY) → \underline{ℳ}_{𝒞}(RX, RZ) relates to the evaluation maps under adjunction—see [D4, Prop. A.3], for instance.

It is now a tedious but routine exercise to prove that the above two maps

RX ⊗ [G\underline{𝒩}_{𝒟} (Y, Z) ⊗ G\underline{𝒩}_{𝒟} (X, Y)] ⇉ RZ

are indeed the same. One forms the adjoints and then writes down a huge commutative diagram. A very similar result (in fact, a special case of the present one) is proven in [D4, A.9].

Note that if G is lax monoidal then it comes with a prescribed map 1_{𝒞} → G(1_{𝒟} ); adjointing gives F (1_{𝒞}) → 1_{𝒟} . The following result concerns compatibility between these maps and γ:

3.7. Proposition. Assume again that G is lax monoidal, and let γ denote a set of four corresponding natural transformations of types (i)–(iv). The following three conditions are equivalent:

(a) For any X, the following square commutes:

\xymatrix{LX \ar[r]^{≅} \ar[d]_{≅} & LX ⊗ 1_{𝒟} \\ L(X ⊗ 1_{𝒞}) \ar[r]^{γ} & LX ⊗ F (1_{𝒞}). \ar[u]}

(b) For any Y, the following square commutes:

\xymatrix{RY \ar[r]^{≅} \ar[d]_{≅} & RY ⊗ 1_{𝒞} \ar[d] \\ R(Y ⊗ 1_{𝒟} ) & RY ⊗ G(1_{𝒟} ). \ar[l]_{γ}}

(c) For any Y, the following square commutes:

\xymatrix{1_{𝒞} \ar[r] \ar[d] & G(1_{𝒟} ) \ar[d] \\ \underline{ℳ}_{𝒞}(RY, RY) & G\underline{𝒩}_{𝒟} (Y, Y) \ar[l]_{γ}}

Proof. Left to the reader.

Finally we have the main definition:

3.8. Definition. Assume given adjoint pairs (F, G) and (L, R) where G is lax monoidal. We will say that (L, R) is an adjoint module over (F, G) if there exists a natural transformation γ: L(X ⊗ c) → LX ⊗ F c such that the conditions of Propositions 3.6(a) and 3.7(a) are both satisfied.

If in addition (F, G) and (L, R) are both Quillen pairs and the equivalent conditions of Proposition 3.5(a) are satisfied we will say that (L, R) is a Quillen adjoint module over (F, G).

3.9. Basic properties. Below we give three properties satisfied by Quillen adjoint modules. Recall the notion of a 𝒞-Quillen adjunction between 𝒞-model categories, as in [D4, A.7]. This is a Quillen pair L: ℳ ⇄ 𝒩: R where ℳ and 𝒩 are 𝒞-model categories, together with natural isomorphisms L(X ⊗ c) ≅ L(X) ⊗ c which reduce to the canonical isomorphism for c = 1_{𝒞} and which are compatible with the associativity isomorphisms in ℳ and 𝒩. See also [Ho1, Def. 4.1.7].

3.10. Proposition. Suppose ℳ and 𝒩 are 𝒞-model categories and L: ℳ ⇄ 𝒩: R is a 𝒞-Quillen adjunction. Then (L, R) is a Quillen adjoint module over the pair (id_{𝒞}, id_{𝒞}).

Proof. Since (L, R) is a 𝒞-adjunction, there are natural isomorphisms LX ⊗ c → L(X ⊗c) which satisfy the associativity and unital properties listed in Propositions 3.6(a) and 3.7(a). This also fulfills the second condition listed in Proposition 3.5(a)

3.11. Proposition. Let F: 𝒞 ⇄ 𝒟: G be a Quillen pair between symmetric monoidal model categories, where G is lax monoidal. Let F′: 𝒟 ⇄ ℰ: G′ be another such pair. Let L: ℳ ⇄ 𝒩: R and L′: 𝒩 ⇄ 𝒫: R′ be Quillen pairs such that (L, R) is a Quillen adjoint module over (F, G) and (L′, R′) is a Quillen adjoint module over (F′, G′). Then (L′L, RR′) is a Quillen adjoint module over (F′F, GG′).

Proof. For X ∈ ℳ and c ∈ 𝒞 we have natural maps

L′L(X ⊗ c) → L′(LX ⊗ F c) → L′(LX) ⊗ F′(F c)

using the adjoint module structure on (L, R) over (F, G) first, and the module structure on (L′, R′) over (F′, G′) second. One just has to check the axioms to see that these maps make (L′L, RR′) a Quillen adjoint module over (F′F, GG′). This is a routine exercise in categorical diagramming which we will leave to the reader.

3.12. Corollary. Suppose (L, R) is a Quillen adjoint module over (F, G), and also suppose that 𝒫 is a 𝒞-model category and J: 𝒫 ⇄ ℳ: K is a 𝒞-Quillen adjunction. Then (LJ, KR) is a Quillen adjoint module over (F, G).

Proof. This is an immediate consequence of the above two propositions.

4. Applications of adjoint modules

Recall from the last section that 𝒞 and 𝒟 are symmetric monoidal model categories, ℳ is a 𝒞-model category, and 𝒩 is a 𝒟-model category. We have Quillen pairs

F: 𝒞 ⇄ 𝒟: G and L: ℳ ⇄ 𝒩: R

in which G is lax monoidal, and we assume that (L, R) is a Quillen adjoint module over (F, G) as defined in Definition 3.8.

Recall the notion of model enrichment from Section 2.3. The assignment X, Y ↦ \underline{𝒩}_{𝒟} (X, Y) is a 𝒟-model enrichment of 𝒩, as in [D4, Example 3.2]. The induced assignment X, Y ↦ G\underline{𝒩}_{𝒟} (X, Y) is a 𝒞-model enrichment of 𝒩, by Proposition 4.6 below. Alternatively, if 𝒬W −↠^{∼} W is a cofibrant-replacement functor for ℳ and W ↣^{∼} ℱW is a fibrant-replacement functor for 𝒩, then one obtains another 𝒞-model enrichment of 𝒩 via X, Y ↦ \underline{ℳ}_{𝒞}(𝒬RℱX, 𝒬RℱY). (This is precisely the enrichment L_{∗}[\underline{ℳ}_{𝒞}], as defined in [D4, Section 3.4].) If R preserves all weak equivalences, the simpler assignment X, Y ↦ \underline{ℳ}_{𝒞}(𝒬RX, 𝒬RY) is also a 𝒞-model enrichment.

4.1. Theorem. Assume the pair (L, R) is a Quillen adjoint module over (F, G). Also assume that G is lax symmetric monoidal and that (L, R) is a Quillen equivalence. Then the two 𝒞-model enrichments on 𝒩 given by X, Y ↦ G\underline{𝒩}_{𝒟} (X, Y) and X, Y ↦ \underline{ℳ}_{𝒞}(𝒬RℱX, 𝒬RℱY) are quasi-equivalent. That is to say, L_{∗}\underline{ℳ}_{𝒞} ≃ G\underline{𝒩}_{𝒟} .

If R preserves all weak equivalences, then the above enrichments are also quasiequivalent to X, Y ↦ \underline{ℳ}_{𝒞}(𝒬RX, 𝒬RY).

We also have the following corollary:

4.2. Corollary. Under the assumptions of the theorem, the two 𝒞-model enrichments on ℳ given by X, Y ↦ G\underline{𝒩}_{𝒟} (ℱL𝒬X, ℱL𝒬Y) and X, Y ↦ \underline{ℳ}_{𝒞}(X, Y) are quasi-equivalent. That is, L^{∗}[G\underline{𝒩}_{𝒟} ] is quasi-equivalent to \underline{ℳ}_{𝒞}. Here 𝒬A −↠^{∼} A and X ↣^{∼} ℱX are now cofibrant- and fibrant- replacement functors in ℳ and 𝒩, respectively.

The quasi-equivalences in Theorem 4.1 and Corollary 4.2 are used in a key argument in [GS] to translate a construction in Hℚ-algebras into rational dgas. The following immediate corollary of the above theorem is what we will mainly need in the present paper.

4.3. Corollary. Assume that 𝒞 is combinatorial, satisfies the monoid axiom, and that 1_{𝒞} is cofibrant. Under the assumptions of the theorem, let X ∈ 𝒩 be a cofibrant-fibrant object. Let A ∈ ℳ be any cofibrant-fibrant object which is weakly equivalent to RX. Then the 𝒞-monoids G\underline{𝒩}_{𝒟} (X, X) and \underline{ℳ}_{𝒞}(A, A) are weakly equivalent.

The extra assumptions on 𝒞 are necessary in order to apply a certain proposition from [D4], saying that quasi-equivalent enrichments give weakly equivalent endomorphism monoids.

The above theorem and corollaries compare enrichments which have been transferred over the right adjoint G. It is more difficult to transfer enrichments over the left adjoint F. The assignment A, B ↦ F \underline{ℳ}_{𝒞}(A, B) is an enrichment only if F is assumed to be strong monoidal; and since \underline{ℳ}_{𝒞}(A, B) will not usually be cofibrant, it will only be a model enrichment under very strong assumptions such as that F preserves all weak equivalences. Finally, the construction F \underline{ℳ}_{𝒞}(A, B) effectively amounts to mixing a left and right adjoint (the mapping object is a right adjoint), and so is not so well-behaved categorically.

The following result is given only for completeness; we have no applications for it at present.

4.4. Proposition. Assume that (L, R) is a Quillen adjoint module over (F, G). Suppose that F and G are both strong monoidal, and that F preserves all weak equivalences. Also assume that G is symmetric monoidal, and that both (F, G) and (L, R) are Quillen equivalences. Then the model enrichments A, B ↦ F \underline{ℳ}_{𝒞}(A, B) and A, B ↦ \underline{𝒩}_{𝒟} (ℱL𝒬A, ℱL𝒬B) are quasi-equivalent.

4.5. Proofs of the above results.

4.6. Proposition. The assignment n, n′ ↦ G\underline{𝒩}_{𝒟} (n, n′) is a 𝒞-model enrichment on 𝒩.

Proof. One uses the monoidal structure on G to produce the associative and unital composition maps. Since G preserves equivalences between all fibrant objects and \underline{𝒩}_{𝒟} (n, n′) is fibrant if n is cofibrant and n′ is fibrant, we see that G\underline{𝒩}_{𝒟} (a′, x) → G\underline{𝒩}_{𝒟} (a, x) and G\underline{𝒩}_{𝒟} (a, x) → G\underline{𝒩}_{𝒟} (a, x′) are weak equivalences whenever a → a′ is a weak equivalence between cofibrant objects and x → x′ is a weak equivalence between fibrant objects.

Proof of Theorem 4.1. For X, Y ∈ 𝒩 define σ(X, Y) = G\underline{𝒩}_{𝒟} (ℱX, ℱY) and τ(X, Y) = \underline{ℳ}_{𝒞}(QRℱX, QRℱY). These are both 𝒞-model enrichments on 𝒩, and the former is quasiequivalent to X, Y ↦ G\underline{𝒩}_{𝒟} (X, Y) by [D4, Prop. 3.9].

Define W (X, Y) = \underline{ℳ}_{𝒞}(𝒬RℱX, RℱY). This is a σ − τ bimodule via the maps

\xymatrix{G\underline{𝒩}_{𝒟} (ℱY, ℱZ) ⊗ \underline{ℳ}_{𝒞}(𝒬RℱX, RℱY) \ar[r]^{γ⊗1} & \underline{ℳ}_{𝒞}(RℱY, RℱZ) ⊗ \underline{ℳ}_{𝒞}(𝒬RℱX, RℱY) \ar[d] \\  & \underline{ℳ}_{𝒞}(𝒬RℱX, RℱZ)}

and

\underline{ℳ}_{𝒞}(𝒬RℱY, RℱZ) ⊗ \underline{ℳ}_{𝒞}(𝒬RℱX, 𝒬RℱY) ⟶ \underline{ℳ}_{𝒞}(𝒬RℱX, RℱZ).

Some routine but tedious checking is required to see that this indeed satisfies the bimodule axioms of [D4, Section 2.2]. This uses the conditions from Proposition 3.6(c) and Proposition 3.7(a).

The canonical maps 𝒬RℱX → RℱX give maps 1_{𝒞} → W (X, X) making W into a pointed bimodule, and one checks using the condition from Proposition 3.5(b) that this is a quasi-equivalence. This last step uses our assumption that (L, R) is a Quillen equivalence.

If R preserves all weak equivalences, then the above proof works even if every appearance of the functor ℱ is removed.

Proof of Corollary 4.2. The result [D4, 3.14(d)] shows that since L is a Quillen equivalence the maps L^{∗} and L_{∗} are inverse bijections. Since we have already proven L_{∗}\underline{ℳ}_{𝒞} ≃ G\underline{𝒩}_{𝒟} , we must have L^{∗}[G\underline{𝒩}_{𝒟} ] ≃ \underline{ℳ}_{𝒞}.

Proof of Corollary 4.3. Using the above theorem together with [D4, Cor. 3.6] (which requires our assumptions on 𝒞) we find that if X ∈ 𝒩 is cofibrant-fibrant then the 𝒞-monoids G\underline{𝒩}_{𝒟} (X, X) and \underline{ℳ}_{𝒞}(𝒬RℱX, 𝒬RℱX) are weakly equivalent. However, note that one has a weak equivalence A ⟶^{∼} 𝒬RℱX. By applying [D4, Cor. 3.7] (in the case where ℐ is the category with one object and an identity map) one finds that the 𝒞-monoids \underline{ℳ}_{𝒞}(𝒬RℱX, 𝒬RℱX) and \underline{ℳ}_{𝒞}(A, A) are weakly equivalent.

Proof of Proposition 4.4. We only sketch this, as the result is not needed in the paper. Consider the following two model enrichments on ℳ:

A, B ↦ \underline{𝒩}_{𝒟} (ℱL𝒬A, ℱL𝒬B)  (4.7)

and

A, B ↦ F G\underline{𝒩}_{𝒟} (ℱL𝒬A, ℱL𝒬B).  (4.8)

The strong monoidal assumptions on F and G show that for any map P ⊗ Q → R in 𝒟, the induced square

\xymatrix{F G(P) ⊗ F G(Q) \ar[r] \ar[d] & F G(R) \ar[d] \\ P ⊗ Q \ar[r] & R}

is commutative. Here the vertical maps come from the counit of the adjunction (F, G), and the top horizontal map uses the isomorphism F G(P ⊗Q) ≅ F G(P)⊗F G(Q) coming from the strong monoidal structures on F and G. This commutative square shows that we have a map of enrichments F G\underline{𝒩}_{𝒟} (X, Y) → \underline{𝒩}_{𝒟} (X, Y) on 𝒩. Using this, we get a map of model enrichments from (4.8) to (4.7), which allows us to think of A, B ↦ \underline{𝒩}_{𝒟} (ℱL𝒬A, ℱL𝒬B) as a bimodule over (4.7) and (4.8) (in either order). Using that (F, G) is a Quillen equivalence one readily checks that this is a quasi-equivalence.

By Corollary 4.2, the model enrichment A, B ↦ G\underline{𝒩}_{𝒟} (ℱL𝒬A, ℱL𝒬B) is quasiequivalent to A, B ↦ \underline{ℳ}_{𝒞}(A, B). But the assumptions on F readily show that appyling F to everything preserves quasi-equivalence. This shows that (4.8) is quasi-equivalent to F \underline{ℳ}_{𝒞}, so (4.7) is also quasi-equivalent to F \underline{ℳ}_{𝒞}. This is what we wanted.

4.9. Applications to module categories. We’ll now apply the above results to the homotopy theory of 𝒞I-categories. Readers may want to review Appendix A before proceeding further.

Let 𝒞 and 𝒟 be cofibrantly-generated symmetric monoidal model categories satisfying the monoid axiom, and assume that 1_{𝒞} and 1_{𝒟} are cofibrant. Let F: 𝒞 ⇄ 𝒟: G be a Quillen pair where G is lax monoidal. Let I be a set and consider the notion of 𝒞Icategory (a category enriched over 𝒞 with object set I) from Appendix A. Note that when I consists of one object then a 𝒞I-category is just a monoid in 𝒞.

Let ℛ be a 𝒟I-category, and consider the category Mod- ℛ of right ℛ-modules. By [SS3, 6.1] the category Mod- ℛ has a model structure in which the weak equivalences and fibrations are obtained by forgetting objectwise to 𝒟. This is a 𝒟-model category in a natural way. The SM7 (or pushout product) condition follows from 𝒟 using [SS1, 3.5] since the 𝒟 action is pointwise, and the unit condition follows from our assumption that 1_{𝒟} is cofibrant (since this implies that the cofibrant ℛ-modules are objectwise cofibrant).

Since G is lax monoidal, Gℛ is a 𝒞I-category and we may consider the corresponding module category Mod- Gℛ. This is a 𝒞-model category. If M is an ℛ-module then GM becomes a Gℛ-module in a natural way, and there is an adjoint pair F^{ℛ} : Mod- Gℛ ⇄ Mod- ℛ: G by Proposition A.6(a). The functors (F^{ℛ} , G) are a Quillen pair since G preserves the objectwise fibrations and trivial fibrations.

We have two Quillen pairs F: 𝒞 ⇄ 𝒟: G and F^{ℛ} : Mod- Gℛ ⇄ Mod- ℛ: G. The categories Mod- Gℛ and Mod- ℛ are 𝒞- and 𝒟-model categories, respectively.

4.10. Proposition. Under the above assumptions on 𝒞, 𝒟, and G one has:

(a) (F^{ℛ} , G) is an adjoint module over (F, G).

(b) If F is strong monoidal, then (F^{ℛ} , G) is a Quillen adjoint module over (F, G).

(c) Assume that 𝒞 is a stable model category whose homotopy category is generated by 1_{𝒞}. Assume as well that F (1_{𝒞}) → 1_{𝒟} is a weak equivalence. Then (F^{ℛ} , G) is a Quillen adjoint module over (F, G).

Proof. In terms of the notation of Section 3 we have L = F^{ℛ} and R = G. A natural transformation γ of the type in Proposition 3.2(iii) is therefore obtained using the lax monoidal structure on G. This automatically satisfies the axioms of Proposition 3.6(b) and Proposition 3.7(b), so that we have an adjoint module over (F, G). This proves (a).

To prove (b) we show that L(X ⊗ c) → LX ⊗ F c is an isomorphism, and hence a weak equivalence. Here L = F^{ℛ} = F (−) ⊗_{F Gℛ} ℛ since F is strong monoidal; see the discussion above [SS3, 3.11]. It is then easy to verify that F^{ℛ} (X ⊗ c) = F (X ⊗ c) ⊗_{F Gℛ} ℛ ≅ (F X ⊗ F c) ⊗_{F Gℛ} ℛ ≅ F^{ℛ} (X) ⊗ F c.

To prove (c), we will verify that G\underline{𝒩}_{𝒟} (LX, Y) ⟶^{γ} \underline{ℳ}_{𝒞}(X, RY) is a weak equivalence whenever X is cofibrant and Y is fibrant. Using our assumption about 1_{𝒞} generating Ho (𝒞), it suffices to show that

[1_{𝒞}, G\underline{𝒩}_{𝒟} (LX, Y)]_{∗} → [1_{𝒞}, \underline{ℳ}_{𝒞}(X, RY)]_{∗}

is an isomorphism of graded groups, where [−, −]_{∗} denotes the graded group of maps in a triangulated category.

By adjointness, the problem reduces to showing that the map

[LX ⊗ F (1_{𝒞}), Y]_{∗} → [L(X ⊗ 1_{𝒞}), Y]_{∗}

is an isomorphism—or in other words, that LX⊗F (1_{𝒞}) → L(X⊗1_{𝒞}) is a weak equivalence. But this follows easily from our assumption that F (1_{𝒞}) → 1_{𝒟} is a weak equivalence.

Now assume that 𝒪 is a cofibrant 𝒞I-category. By Proposition A.3 there is an adjunction F^{𝒟I} : 𝒞I − 𝒞at ⇄ 𝒟I − 𝒞at: G, so that we get a 𝒟I-category F^{𝒟I} 𝒪. By Proposition A.6(b) there is a Quillen pair

F_{𝒪} : Mod- 𝒪 ⇄ Mod- F^{𝒟I} 𝒪: G_{𝒪} .

4.11. Proposition. In the above setting one has:

(a) (F_{𝒪} , G_{𝒪}) is an adjoint module over (F, G).

(b) If F is strong monoidal, then (F_{𝒪} , G_{𝒪}) is a Quillen adjoint module over (F, G).

(c) Assume that 𝒞 is a stable model category whose homotopy category is generated by 1_{𝒞}, and that F (1_{𝒞}) → 1_{𝒟} is a weak equivalence. Then (F_{𝒪} , G_{𝒪}) is a Quillen adjoint module over (F, G).

Proof. Write ℛ = F^{𝒟I} 𝒪. The adjunction (F_{𝒪} , G_{𝒪}) is the composite of the two adjunctions

\xymatrix{Mod- 𝒪 \ar[r]^{β_{∗}} & Mod- Gℛ \ar[r]^{F^{ℛ}} \ar[l]^{β^{∗}} & Mod- ℛ \ar[l]^{G}}

where β: 𝒪 → Gℛ = GF^{𝒟I} 𝒪 is the unit of the adjunction (F^{𝒟I} , G).

But (β_{∗}, β^{∗}) is a 𝒞-Quillen adjunction and by Proposition 4.10, under either set of conditions, we know (F^{ℛ} , G) is a Quillen adjoint module over (F, G). The result now follows immediately from Corollary 3.12.

4.12. Corollary. In addition to our previous assumptions, assume that G is lax symmetric monoidal and 𝒪 is a cofibrant 𝒞I-category. Suppose also that (F_{𝒪}, G_{𝒪}) is a Quillen equivalence and the hypotheses in either part (b) or (c) hold from Proposition 4.11. Let X ∈ Mod-(F^{𝒟I} 𝒪) be a cofibrant-fibrant object and let A ∈ Mod- 𝒪 be any module weakly equivalent to G_{𝒪} X. Then the 𝒞-monoids

\underline{Mod- 𝒪}_{𝒞}(A, A)  and  G[\underline{Mod-(F^{𝒟I} 𝒪)}_{𝒟} (X, X)]

are weakly equivalent.

Proof. This follows from the above proposition and Corollary 4.3.

4.13. Example. The adjoint pair L: Sp^{Σ}(ch_{+}) ⇄ Sp^{Σ}(s𝒜b): ν from [S, 4.3] forms one example for (F, G). The result [S, 3.4] shows that Sp^{Σ}(ch_{+}) and Sp^{Σ}(s𝒜b) are cofibrantly generated symmetric monoidal model categories which satisfy the monoid axiom. The conditions in Proposition 4.10(c) or 4.11(c) are verified in the last paragraph of the proof of [S, 4.3]. Note, though, that L is not strong monoidal. This failure is due to the fact that the adjunction N: s𝒜b ⇄ ch_{+} : Γ is not monoidal [SS3, 2.14].

Corollary 4.12 holds for (L, ν) in place of (F, G) because N is lax symmetric monoidal, so its prolongation and ν are also lax symmetric monoidal. The fact that (L_{𝒪} , (ν)_{𝒪}) is a Quillen equivalence follows from [S, 3.4, 4.3] and [SS3, 6.5(1)]. See also Proposition A.6(c).

5. Transporting enrichments

In this section we prove a technical result about transporting enrichments. This will be needed later, in the proof of Proposition 9.4. The basic idea is as follows. Suppose ℳ is a 𝒞-model category, where 𝒞 is a certain symmetric monoidal model category. Assume also that 𝒟 is another symmteric monoidal model category, and that one has a Quillen equivalence 𝒞 ⇄ 𝒟 which is compatible with the monoidal structure. Then one might hope to find a 𝒟-model category 𝒩 which is Quillen equivalent to ℳ, and where the Quillen equivalence aligns the 𝒞- and 𝒟-structures. In this section we prove one theorem along these lines, assuming several hypotheses on the given data.

We begin with the following two definitions:

5.1. Definition. Let 𝒯 be a triangulated category with infinite coproducts.

(a) An object P ∈ 𝒯 is called compact if ⊕_{α}𝒯(P, X_{α}) → 𝒯(P, ⊕_{α}X_{α}) is an isomorphism for every set of objects {X_{α}};

(b) A set of objects S ⊆ 𝒯 is a generating set if the only full, triangulated subcategory of 𝒯 which contains S and is closed under arbitrary coproducts is 𝒯 itself. If S is a singleton set {P } we say that P is a generator.

When ℳ is a stable model category we will call an object compact if it is compact in ℋo(ℳ), and similarly for the notion of generating set. Most stable model categories of interest have a generating set of compact objects. For example, Hovey shows in [Ho1, 7.4.4] that this is true for any finitely-generated, stable model category.

Let 𝒞 and 𝒟 be symmetric monoidal, stable model categories. Let ℳ be a pointed 𝒞-model category (so that ℳ is also stable). We make the following assumptions:

(a) 𝒞 and 𝒟 are combinatorial model categories satisfying the monoid axiom, and their units are cofibrant.

(b) There is a weak monoidal Quillen equivalence F: 𝒞 ⇄ 𝒟: G, where G is lax symmetric monoidal.

(c) 𝒞 satisfies axioms (QI1-2) from Appendix A.

(d) 𝒞 is a stable model category whose homotopy category is generated by 1_{𝒞}, and F (1_{𝒞}) → 1_{𝒟} is a weak equivalence.

(e) ℳ has a generating set of compact objects.

If 𝒩 is a 𝒟-model category, let G\underline{𝒩}_{𝒟} denote the assignment X, Y ↦ G\underline{𝒩}_{𝒟} (X, Y). By Proposition 4.6 this is a 𝒞-model enrichment of 𝒩.

5.2. Proposition. Under the above conditions there exists a combinatorial 𝒟-model category 𝒩 and a zig-zag of Quillen equivalences

ℳ ⟵^{L_{1}} ℳ_{1} ⟶^{L_{2}} 𝒩

such that the model enrichment \underline{ℳ}_{𝒞} is quasi-equivalent to (L_{1})_{∗}(L_{2})^{∗}[G\underline{𝒩}_{𝒟} ].

If, in addition, 𝒞 and 𝒟 are additive model categories (see the following section for the definition) then ℳ_{1} and 𝒩 may also be chosen to be additive.

By [D4, 3.6] this yields the following immediate corollary:

5.3. Corollary. If Y ∈ 𝒩 and X ∈ ℳ are cofibrant-fibrant objects and Y is the image of X under the derived functors of the above Quillen equivalence ℳ ≃ 𝒩, then the 𝒞-monoids \underline{ℳ}_{𝒞}(X, X) and G\underline{𝒩}_{𝒟} (Y, Y) are weakly equivalent.

Proof of Proposition 5.2. Constructing the model category 𝒩 will require several steps, and we will start by just giving a sketch—then we will come back and provide detailed justifications afterwards.

Let I denote a set of cofibrant-fibrant, compact objects which generate ℳ. Let 𝒪 be the 𝒞I-category [Bo, 6.2] defined by 𝒪(i, j) = \underline{ℳ}_{𝒞}(i, j). Then there is a 𝒞-Quillen equivalence

T: Mod- 𝒪 ⇄ ℳ: S  (5.4)

where Mod- 𝒪 is the model category of right 𝒪-modules (see Proposition A.2).

Let g: \overline{𝒪} → 𝒪 be a cofibrant-replacement for 𝒪 in the model category of 𝒞I-categories (Proposition A.3(a)). Then tensoring and restricting give the left and right adjoints of a 𝒞-Quillen equivalence

g_{∗}: Mod- \overline{𝒪} ⇄ Mod- 𝒪: g^{∗}  (5.5)

(see Proposition A.2(c)).

Next we use the functor L^{𝒟I} from Proposition A.3(b). This gives us a 𝒟I-category L^{𝒟I} \overline{𝒪} and a Quillen equivalence

L\overline{𝒪} : Mod- \overline{𝒪} ⇄ Mod-(L^{𝒟I} \overline{𝒪}): ν.  (5.6)

Let 𝒩 = Mod-(L^{𝒟I} \overline{𝒪}). This is a 𝒟-model category, and we have established a zig-zag of Quillen equivalences

ℳ ⟵^{∼} Mod- 𝒪 ⟵^{∼} Mod- \overline{𝒪} ⟶^{∼} Mod-(L^{𝒟I} \overline{𝒪}) = 𝒩.

We set ℳ_{1} = Mod- \overline{𝒪}. Note that if 𝒞 and 𝒟 are additive model categories then by Corollary 6.9 so are ℳ_{1} and 𝒩 (since ℳ_{1} is a 𝒞-model category and 𝒩 is a 𝒟-model category).

Now we fill in the details of the above sketch. The category of right 𝒪-modules Mod- 𝒪 is defined in [SS3, Section 6], and the model structure on Mod- 𝒪 is provided in [SS3, 6.1(1)]. See Appendix A for a review. To justify the Quillen equivalence in (5.4), define S: ℳ → Mod- 𝒪 by letting S(Z) be the functor i ↦ Hom_{𝒞}(i, Z). This obviously comes equipped with a structure of right 𝒪-module. The construction of the left adjoint can be copied almost verbatim from [SS2, 3.9.3(i)], which handled the case where 𝒞 was Sp^{Σ}. The right adjoint obviously preserves fibrations and trivial fibrations, so we have a Quillen pair. It is readily seen to be a 𝒞-Quillen pair.

Finally, that this is a Quillen equivalence follows just as in [SS2, 3.9.3(ii)]; this uses that I was a generating set of compact objects. The proof can be summarized quickly as follows. First, the compactness of the objects in I shows that the derived functor of S preserves all coproducts; this is trivially true for the derived functor of T because it is a left adjoint. One has canonical generators F r_{i} ∈ Mod- 𝒪 for each i ∈ I, and adjointness shows that T (F r_{i}) ≅ i. Likewise, S(i) ≅ F r_{i}. Using that the derived functors of S and T preserve coproducts and triangles, one now deduces that the respective composites are naturally isomorphic to the identities. This completes step (5.4) above.

We now turn to (5.5). The map of 𝒞I-categories g: \overline{𝒪} → 𝒪 gives a Quillen map Mod- \overline{𝒪} → Mod- 𝒪 by Proposition A.2(b). We will know this is a Quillen equivalence by Proposition A.2(c) as long as we know that 𝒞 satisfies the axioms (QI1-2) of Appendix A.

The Quillen equivalence of (5.6) is a direct application of Proposition A.6(c).

At this point we have constructed the zig-zag ℳ ⟵^{L_{1}} ℳ_{1} ⟶^{L_{2}} 𝒩. We must verify that (L_{1})_{∗}(L_{2})^{∗}[G\underline{𝒩}_{𝒟} ] is quasi-equivalent to \underline{ℳ}_{𝒞}.

It follows from Proposition 4.11 and Theorem 4.1 that (L_{2})^{∗}[G\underline{𝒩}_{𝒟} ] is quasi-equivalent to (\underline{ℳ}_{1})_{𝒞}. This is where the theory of adjoint modules was needed. Since L_{1} is a 𝒞-Quillen equivalence, it follows from [D4, 3.14(e)] that (L_{1})_{∗}[(\underline{ℳ}_{1})_{𝒞}] is quasi-equivalent to \underline{ℳ}_{𝒞}. So these two statements give exactly what we want.

6. Additive model categories

Now the second half of the paper begins. We change direction and start to pursue our main results on additive enrichments. In the present section we define the notion of an additive model category, and prove some basic results for recognizing them.

A category is preadditive if its hom-sets have natural structures of abelian groups for which the composition pairing is biadditive. A category is additive if it is preadditive and it has finite coproducts. This forces the existence of an initial object (the empty coproduct), which will necessarily be a zero object. See [ML, Section VIII.2]. A functor F: 𝒞 → 𝒟 between additive categories is an additive functor if F (f +g) = F (f)+F (g) for any two maps f, g: X → Y.

Now let ℳ be a model category whose underlying category is additive. Write ℳ_{cof} for the full subcategory of cofibrant objects, and cℳ for the category of cosimplicial objects in ℳ. Recall from [Hi, Section 15.3] that cℳ has a Reedy model category structure. Also recall that a cosimplicial resolution is a Reedy cofibrant object of cℳ in which every coface and codegeneracy map is a weak equivalence.

6.1. Definition. Let ℐ be a small, additive subcategory of ℳ_{cof}. By an additive cosimplicial resolution on ℐ we mean an additive functor Γ: ℐ → cℳ whose image lies in the subcategory of cosimplicial resolutions, together with a natural weak equivalence

Γ(X)^{0} ⟶^{∼} X.

By [Hi, 16.1.9], any small subcategory ℐ ⊆ ℳ_{cof} has a cosimplicial resolution; however, the existence of an additive cosimplicial resolution is not at all clear.

If Γ and Γ′ are two additive cosimplicial resolutions on ℐ, then define a map Γ → Γ′ to be a natural transformation of functors which gives commutative triangles

\xymatrix{Γ(X)^{0} \ar[d] \ar[r] & X \\ Γ′(X)^{0} \ar[ur]}

for all X ∈ ℐ. The map is called a weak equivalence if all the maps Γ(X) → Γ′(X) are weak equivalences.

6.2. Definition. A model category ℳ is additive if its underlying category is additive and if for every small, full subcategory ℐ of ℳ_{cof} the following two statements are satisfied:

(a) ℐ has an additive cosimplicial resolution;

(b) The category of additive cosimplicial resolutions on ℐ, where maps are natural weak equivalences, has a contractible nerve.

6.3. Proposition. Let ℳ be a model category whose underlying category is additive. Suppose that there is a functor F: ℳ_{cof} → cℳ together with a natural isomorphism F^{0}(X) ≅ X. Assume that each F (X) is a cosimplicial resolution, that F preserves colimits, and that if X ↣ Y is a cofibration then F (X) → F (Y) is a Reedy cofibration. Then ℳ is an additive model category.

Note that the functor F will automatically be additive; since it preserves colimits, it preserves direct sums.

Proof. The existence of additive cosimplicial resolutions is provided by F. So we must only prove the contractibility of the category of all such resolutions.

If Γ ∈ cℳ is any cosimplicial object, applying F to Γ yields a bi-cosimplicial object F Γ given by [m], [n] ↦ F^{m}Γ^{n}. Let Γ̃ ∈ cℳ denote the diagonal of this bi-cosimplicial object, and note that there is a natural map Γ̃ → Γ. We claim that if Γ is a cosimplicial resolution then so is Γ̃.

Suppose that Γ ∈ cℳ is a cosimplicial resolution of some object X. Then every latching map L^{n}Γ → Γ^{n} is a cofibration (see [Hi, 15.3] for a discussion of latching maps). From the bi-cosimplicial object F Γ, we get a ‘vertical’ latching map in cℳ of the form L^{∗,n}[F Γ] → F (Γ^{n}). Here the domain is the cosimplicial object which in level m is the nth latching object of [F Γ]^{m,∗}. Since the latching spaces are formed as colimits, and F preserves colimits, one has L^{∗,n}[F Γ] ≅ F (L^{n}Γ). So our vertical latching map is F (L^{n}Γ) → F (Γ^{n}). But this is the result of applying F to a cofibration in ℳ, so it is a Reedy cofibration.

So we are in the situation of Lemma 6.5 below, in which every vertical latching map of F Γ is a Reedy cofibration. By the lemma, this implies that the diagonal Γ̃ is Reedy cofibrant. Since clearly every map in F Γ is a weak equivalence, it is therefore a cosimplicial resolution of X.

If Γ is a cosimplicial resolution of X, then the weak equivalence Γ(X)^{0} ⟶^{∼} X gives a map of cosimplicial objects Γ → cX (here cX denotes the constant cosimplicial object with X in every dimension). Applying ˜(−), we get a diagram of cosimplicial objects

\xymatrix{Γ̃ \ar[r] \ar[d] & Γ \ar[d] \\ { \atop (cX)} \ar[r] & cX.}  (6.4)

Note that { \atop (cX)} is just F (X), which is a cosimplicial resolution by the assumptions of the proposition. All the maps in the above square can be checked to be levelwise weak equivalences.

Now suppose that ℐ is a small, full subcategory of ℳ_{cof}. Let CR(ℐ) be the category of all additive cosimplicial resolutions on ℐ, where the maps are natural weak equivalences. The restriction of F to a functor ℐ → cℳ is an object of CR(ℐ), which we also denote F by abuse. Consider the overcategory (CR(ℐ) ↓ F) together with the forgetful functor U: (CR(ℐ) ↓ F) → CR(ℐ) sending an object [Γ → F] to Γ.

We define a functor CR(ℐ) → (CR(ℐ) ↓ F) by Γ ↦ Γ̃ (we are applying the construction ˜(−) to every Γ(X), for X ∈ ℐ). The cosimplicial resolution Γ̃ is augmented over F by the left vertical arrow in (6.4).

Consider the two functors

CR(ℐ) ⟶ (CR(ℐ) ↓ F) ⟶^{U} CR(ℐ).

Then there is a natural transformation from the composite to the identity, via the natural maps { \atop Γ(X)} → Γ(X). This shows that on the level of nerves, the identity map for CR(ℐ) is homotopic to a map which factors through a contractible space—the nerve of (CR(ℐ) ↓ F) is contractible because this category has a final object. So we conclude that CR(ℐ) has contractible nerve.

We need some notation for the following lemma. Let X^{∗,∗} be a bi-cosimplicial object in a model category ℳ. Considering this as an object of c(cℳ), one obtains a ‘vertical’ latching map L^{∗,n}X → X^{∗,n} in cℳ. Here L^{∗,n}X denotes the cosimplicial object sending [m] to the nth latching object of X^{m,∗}.

6.5. Lemma. Let ℳ be any model category. Suppose that X^{∗,∗} is a bi-cosimplicial object of ℳ—that is, X ∈ c(cℳ). Assume that every latching map L^{∗,n}X → X^{∗,n} is a Reedy cofibration in cℳ. Then the diagonal cosimplicial object [n] ↦ X^{n,n} is Reedy cofibrant.

The proof of the above lemma is a little technical. We defer it until the end of the section.

6.6. Corollary. Let 𝒞 and ℳ be model categories, where the underlying category of ℳ is additive. Suppose there is a bifunctor ⊗: ℳ × 𝒞 → ℳ satisfying the pushout-product axiom for cofibrations: if i: A ↣ B is a cofibration in ℳ and j: X ↣ Y is a cofibration in 𝒞, then (A ⊗ Y) ⨿_{(A⊗X)} (B ⊗ X) → B ⊗ Y is a cofibration which is a weak equivalence if either i of j is. Suppose also that

(i) For any X ∈ 𝒞 the functor (−) ⊗ X preserves colimits;

(ii) For any A ∈ ℳ the functor A ⊗ (−) preserves colimits;

(iii) There is a cofibrant object 1 ∈ 𝒞 and natural isomorphisms A ⊗ 1 ≅ A.

Then ℳ is an additive model category.

Proof. Let Γ ∈ c𝒞 be a cosimplicial resolution of 1 with Γ^{0} = 1. For any cofibrant object A ∈ ℳ, let F (A) be the cosimplicial object [n] ↦ A ⊗ Γ^{n}. The pushout-product axiom, together with assumption (ii), shows that F (A) is a cosimplicial resolution of A. Assumption (i) implies that F preserves colimits, and assumption (iii) says there are natural isomorphisms F (A)^{0} ≅ A. Finally, it is an easy exercise to use assumption (ii) and the pushout-product axiom to show that if A → B is a cofibration then F (A) → F (B) is a Reedy cofibration. The result now follows by applying Proposition 6.3.

The above corollary lets one identify many examples of additive model categories. We only take note of the few obvious ones:

6.7. Corollary. If R is a ring, consider the model category s(R−ℳod) where fibrations and weak equivalences are determined by the forgetful functor to s𝒮et. This is an additive model category. So is the model category Ch(R) of unbounded chain complexes, where weak equivalences are quasi-isomorphisms and fibrations are sujections.

Proof. This results from two applications of the previous corollary. For the first statement we take ℳ = s(R−ℳod), 𝒞 = s(ℤ−ℳod), and ⊗ to be the levelwise tensor product over ℤ. Here we are using that if M is an R-module and A is a ℤ-module then M ⊗_{ℤ} A has a natural R-module structure from the left.

For the second statement we can take ℳ = Ch(R), 𝒞 = Ch_{≥0}(ℤ), and ⊗ the usual tensor product of chain complexes over ℤ. (One could also take 𝒞 = Ch(ℤ), but verifying the pushout-product axiom is a little easier for bounded below complexes).

If R is a dga, then R − ℳod has a model category structure where weak equivalences are quasi-isomorphisms and fibrations are surjections.

6.8. Corollary. If R is a dga, then the model category R − ℳod is additive.

Proof. We again apply Corollary 6.6, this time with ℳ = R − ℳod and 𝒞 = Ch_{≥0}(ℤ). The ⊗ functor is the tensor product M, C ↦ M ⊗_{ℤ} C with the induced left R-module structure.

We also note the following result:

6.9. Corollary. Let 𝒞 be a symmetric monoidal model category in which the unit is cofibrant, and where the underlying category is additive. Then 𝒞 is an additive model category. Any 𝒞-model category is also additive.

Proof. The first statement follows immediately from Corollary 6.6, as the bifunctor X, Z ↦ X ⊗ Z preserves colimits in both variables.

The second statement is also a direct application of Corollary 6.6, as soon as one notes that if ℳ is a 𝒞-model category then the underlying category of ℳ is additive. This follows using the adjunctions ℳ(X, Y) ≅ ℳ(X ⊗ 1_{𝒞}, Y) ≅ 𝒞(1_{𝒞}, \underline{ℳ}_{𝒞}(X, Y)), as there is a natural abelian group structure on the latter set. One checks that composition is biadditive with respect to this structure.

6.10. Bisimplicial machinery. The last thing we must do in this section is prove Lemma 6.5. This requires some machinery which we briefly recall.

If K ∈ s𝒮et and A ∈ cℳ, one may form the coend A⊗K ∈ ℳ. This is the coequalizer of the two arrows

∐_{[n]↦[m]} A^{n} ⊗ K_{m} ⇉ ∐_{[n]} A^{n} ⊗ K_{n}  (6.11)

where A^{n} ⊗K_{m} is shorthand for a coproduct of copies of A^{n} indexed by the set K_{m}. There are adjunctions

ℳ(A ⊗ K, X) ≅ cℳ(A, X^{K}) ≅ s𝒮et(K, ℳ(A, X))  (6.12)

for A ∈ cℳ, K ∈ s𝒮et, and X ∈ ℳ. Here X^{K} is the cosimplicial object [n] ↦ X^{K_{n}}, where X^{K_{n}} denotes a product of copies of X indexed by the set K_{n}. One checks—using the above adjunctions or otherwise—that A ⊗ Δ^{n} ≅ A^{n}, and A ⊗ ∂Δ^{n} is isomorphic to the nth latching object of A [Hi, Def. 15.2.5]. See [D1, Section 4] for the dual situation with sℳ instead of cℳ.

Write s^{2}𝒮et for the category of bisimplicial sets and c^{2}ℳ for the category of bicosimplicial objects in ℳ. When drawing a bisimplicial set P we will draw each P_{m,∗} horizontally, and each P_{∗,n} vertically. If K ∈ s𝒮et and P ∈ s^{2}𝒮et, let vMap(K, P) denote the simplicial set [n] ↦ s𝒮et(K, P_{∗,n}). We are mapping K into the vertical simplicial sets of P.

If K, L ∈ s𝒮et write K ⊠ L for the bisimplicial set [m], [n] ↦ K_{m} × L_{n}. Observe that there is an adjunction formula

s^{2}𝒮et(K ⊠ L, P) ≅ s𝒮et(L, vMap(K, P)).  (6.13)

Note in particular that s^{2}𝒮et(Δ^{m} ⊠ Δ^{n}, P) ≅ P_{m,n}.

If P ∈ s^{2}𝒮et and A ∈ c^{2}ℳ, one can form a coend A ⊗ P ∈ ℳ similarly to what was done in (6.11). There are adjunction formulas analogous to (6.12). One checks that A ⊗ (Δ^{m} ⊠ Δ^{n}) ≅ A^{m,n}, and more generally A ⊗ (Δ^{m} ⊠ L) ≅ A^{m,∗} ⊗ L (use (6.13) for both). So, for instance, A ⊗ (Δ^{m} ⊠ ∂Δ^{n}) is the nth latching object for the cosimplicial object A^{m,∗}.

Finally, recall from [BF, p. 125] that the diagonal functor diag: s^{2}𝒮et → s𝒮et has a left adjoint which we will call d: s𝒮et → s^{2}𝒮et. It follows immediately from adjointness that dΔ^{n} ≅ Δ^{n} ⊠ Δ^{n}. Since d preserves colimits and every simplicial set is a colimit of Δ^{n}’s, this tells us what d does to any simpiclial set.

By chasing through adjunctions one finds that if X ∈ c^{2}ℳ and K ∈ s𝒮et then diag(X) ⊗ K ≅ X ⊗ dK.

Proof of Lemma 6.5. Consider the object X ∈ c^{2}ℳ given in the statement of the lemma. Our task is to show that diag(X) ⊗ ∂Δ^{n} → diag(X) ⊗ Δ^{n} is a cofibration, for each n. This is the condition for diag(X) to be Reedy cofibrant. Using the isomorphisms diag(X) ⊗ K ≅ X ⊗ dK, this is equivalent to showing that the map X ⊗ d(∂Δ^{n}) → X ⊗ dΔ^{n} ≅ X ⊗ (Δ^{n} ⊠ Δ^{n}) is a cofibration.

Let S denote the set of all maps P → Q of bisimplicial sets such that X ⊗ P → X ⊗ Q is a cofibration in ℳ. This set is closed under composition and cobase change. Our assumption about the latching maps of X amounts to saying that the maps

(∂Δ^{k} ⊠ Δ^{n}) ⨿_{(∂Δ^{k}⊠∂Δ^{n})} (Δ^{k} ⊠ ∂Δ^{n}) ⟶ Δ^{k} ⊠ Δ^{n}  (6.14)

belong to S, for all n and k. Our goal is to show that this forces d(∂Δ^{n}) → Δ^{n} ⊠ Δ^{n} to also belong to S.

We have now reduced things to a problem in combinatorial homotopy theory. Namely, we must show that d(∂Δ^{n}) → Δ^{n}⊠Δ^{n} can be obtained from the maps in (6.14) by iterated cobase changes and compositions. But a little thought shows that every monomorphism of bisimplicial sets can be obtained in this way from the maps in (6.14) (the point is that every monomorphism of simplicial sets can be obtained from the maps ∂Δ^{n} → Δ^{n} in the same way). So we are done.

7. Universal additive model categories

Suppose 𝒞 is a small category. The paper [D2] introduced the idea of a universal model category built from 𝒞, there denoted U𝒞. This is just the category of functors Func(𝒞^{op}, s𝒮et) with a well-known model structure. If 𝒞 is also an additive category then one can ask for a universal additive model category built from 𝒞. This section develops this concept.

7.1. Presheaves and additive presheaves. Let 𝒞 be a small, additive category. Let Func(𝒞^{op}, 𝒜b) denote the category of all functors. Note that for every X ∈ 𝒞, the representable functor rX: 𝒞^{op} → 𝒜b defined by U ↦ 𝒞(U, X) is additive.

The Yoneda Lemma does not hold in Func(𝒞^{op}, 𝒜b): that is, if F ∈ Func(𝒞^{op}, 𝒜b) one need not have Hom(rX, F) ≅ F (X) for all X ∈ 𝒞. But it is easy to check that this does hold when F is an additive functor.

Let Func_{ad}(𝒞^{op}, 𝒜b) denote the full subcategory of additive functors. The following lemma records several basic facts about this category.

7.2. Lemma. Let 𝒞 be a small, additive category.

(a) Colimits and limits in Func_{ad}(𝒞^{op}, 𝒜b) are the same as those in Func(𝒞^{op}, 𝒜b).

(b) Every additive functor F ∈ Func(𝒞^{op}, 𝒜b) is isomorphic to its canonical colimit with respect to the embedding r: 𝒞 ↪ Func(𝒞^{op}, 𝒜b). That is, the natural map [colim_{rX→F}(rX)] → F is an isomorphism.

(c) The additive functors in Func(𝒞^{op}, 𝒜b) are precisely those functors which are colimits of representables.

(d) The inclusion i: Func_{ad}(𝒞^{op}, 𝒜b) ↪ Func(𝒞^{op}, 𝒜b) has a left adjoint Ad (for ‘additivization’), and the composite Ad ∘i is naturally isomorphic to the identity.

(e) Suppose given a co-complete, additive category 𝒜 and an additive functor γ: 𝒞 → 𝒜. Define Sing: 𝒜 → Func_{ad}(𝒞^{op}, 𝒜b) by letting Sing(a) be the functor c ↦ 𝒜(γc, a). Then Sing has a left adjoint Re, and there are natural isomorphisms Re(rX) ≅ γ(X).

Proof. We mostly leave this to the reader. We note, however, that the fact that 𝒞 has finite coproducts (which is part of the definition of an additive category) is needed in (b). This ensures that the categories indexing the canonical colimits are pseudo-filtered, in the sense that for any objects i and j there is a third object k and maps i → k, j → k. The proof uses the fact that if D: I → 𝒜b, where I is pseudo-filtered, then for every element x ∈ colim_{I} D there exists an i ∈ I such that x is in the image of D_{i} → colim_{I} D.

Also, we define the additivization functor from (d). If F is any functor, then (Ad F)(X) is the quotient of F (X) by the subgroup generated by all (f + g)^{∗}(s) − f^{∗}(s) − g^{∗}(s) for all objects Y, all functions f, g: X → Y, and all s ∈ F (Y).

For the proof of (e), note the following. If X ∈ 𝒜 and B is an abelian group, one can define X ⊗ B as the coequalizer of two maps

∐_{B×B} X ⇉ ∐_{B} X.

Here the objects are coproducts of copies of X, indexed by the sets B × B and B, respectively. To describe the two maps, we have to say what they do to each summand corresponding to a pair (b_{1}, b_{2}). The first map is just the inclusion into the summand indexed by b_{1} + b_{2}. The second map is the sum of the two inclusion maps corresponding to the summands b_{1} and b_{2}. One checks that with this definition there is a natural adjunction isomorphism 𝒜(X ⊗ B, Y) ≅ 𝒜b(B, 𝒜(X, Y)).

Recall that we are given an additive functor γ: 𝒞 → 𝒜. Given a functor F: 𝒞^{op} → 𝒜b, we consider the coend

γ ⊗ F = coeq[∐_{c→d} γ(c) ⊗ F (d) ⇉ ∐_{c} γ(c) ⊗ F (c)].

When F is an additive functor one defines Re(F) = γ ⊗ F. It is routine to check that this is a left adjoint to Sing.

By [Hi, Th. 11.6.1] the category Func(𝒞^{op}, s𝒜b) has a cofibrantly-generated model structure in which the weak equivalences and fibrations are defined objectwise. We will need the analogous result for the category of additive functors:

7.3. Lemma. Let 𝒞 be a small, preadditive category. Then Func_{ad}(𝒞^{op}, s𝒜b) has a cofibrantly-generated model structure in which the weak equivalences and fibrations are defined objectwise. This model structure is simplicial, left proper, combinatorial, and cellular.

Proof. The proof uses the adjoint pair (Ad, i) to create the model structure, as in [Hi, Th. 11.3.2]. Recall that the model category Func(𝒞^{op}, s𝒜b) has generating trivial cofibrations J = {rX × ℤ[Λ^{n,k}] → rX × ℤ[Δ^{n}] | X ∈ 𝒞}. Our notation is that if K ∈ s𝒮et then ℤ[K] ∈ s𝒜b is the levelwise free abelian group on K; and if A ∈ s𝒜b then rX ×A denotes the presheaf U ↦ 𝒞(U, X) × A (with the product performed levelwise). Note that we think of rX as a 𝒮et-valued functor here, so 𝒞(U, X) × A denotes a direct sum of copies of A indexed by the set 𝒞(U, X)—this is not the same as as the direct product of the abelian groups 𝒞(U, X) and A.

To apply [Hi, 11.3.2] we must verify that the functor i takes relative Ad(J)-cell complexes to weak equivalences. However, note that if A is an abelian group then Ad(rX × A) ≅ rX ⊗ A, where the latter refers to the presheaf U ↦ 𝒞(U, X) ⊗ A. So Ad(J) is the set of maps rX ⊗ ℤ[Λ^{n,k}] → rX ⊗ ℤ[Δ^{n}]. Objectwise, these maps are monomorphisms and weak equivalences of simplicial abelian groups.

Now, the model category s𝒜b has the special property that a pushout of a map which is both a monomorphism and a weak equivalence is still a monomorphism and weak equivalence. The fact that forming pushouts in Func_{ad}(𝒞^{op}, s𝒜b) and Func(𝒞^{op}, s𝒜b) give the same answers (by Lemma 7.2(a)) and are done objectwise therefore shows that the Ad(J)-cell complexes are objectwise monomorphisms and objectwise weak equivalences. In particular, they are weak equivalences in Func(𝒞^{op}, s𝒜b).

Finally, it is routine to check that the resulting model structure is simplicial, left proper, combinatorial, and cellular.

From now on we will write U_{ad}𝒞 for the category Func_{ad}(𝒞^{op}, s𝒜b) with the model structure provided by the above lemma. The reason for the notation is provided by the next result.

Recall that if L_{1}, L_{2} : ℳ → 𝒩 are two Quillen maps then a Quillen homotopy from L_{1} to L_{2} is a natural transformation L_{1} → L_{2} which is a weak equivalence on the cofibrant objects.

If ℳ is a model category and S is a set of maps in ℳ, then we use ℳ/S to denote the left Bousfield localization of ℳ at S, if it exists. See [Hi, Chapters 3–4] and [D2] for a discussion. The localizations always exist when ℳ = U_{ad}𝒞, since this model category is left proper and cellular.

7.4. Theorem. Let ℳ be an additive model category.

(a) Let 𝒞 be a small, additive category and γ: 𝒞 → ℳ an additive functor taking values in the cofibrant objects. Then there is a Quillen pair Re: U_{ad}𝒞 ⇄ ℳ: Sing together with a natural weak equivalence η: Re ∘r ⟶^{∼} γ. The category whose objects are such triples (Re, Sing, η) and whose maps are Quillen homotopies compatible with η—defined just as in [D2, paragraph preceding Proposition 1.1]—is contractible.

(b) If ℳ is combinatorial then there is a Quillen equivalence U_{ad}𝒞/S ⟶^{∼} ℳ for some small, additive category 𝒞 and some set of maps S in U_{ad}𝒞.

(c) Suppose ℳ ⟵^{∼} ℳ_{1} ⟶^{∼} ⋯ ⟵^{∼} ℳ_{n} ⟶^{∼} 𝒩 is a zig-zag of Quillen equivalences in which all the model categories are additive. If ℳ is combinatorial, there is a simple zig-zag of equivalences

ℳ ⟵^{∼} U_{ad}𝒞/S ⟶^{∼} 𝒩

such that the derived equivalence Ho (ℳ) ≃ Ho (𝒩) is isomorphic to the derived equivalence given by the original zig-zag.

Proof. For (a), one shows that giving a Quillen pair Re: U_{ad}𝒞 ⇄ ℳ: Sing together with a natural weak equivalence Re(rX) ⟶^{∼} γ(X) is precisely the same as giving an additive cosimplicial resolution on γ. The proof of this is exactly the same as [D2, Prop. 3.4]. Giving a Quillen homotopy between two such Quillen pairs exactly amounts to giving a natural weak equivalence between the corresponding cosimplicial resolutions. This proves (a), once one recalls our definition of additive model categories.

The proof for (c) now exactly follow the case for U𝒞 given in [D2, Cor. 6.5]. One uses along the way that adjoint functors between additive categories are necessarily additive functors.

The proof of (b) is slightly more complicated; we will return to it at the end of this section, after some discussion.

7.5. Remark. Part of the result in (a) states that the category whose objects are triples (Re, Sing, η) is contractible. This is what distinguishes U_{ad}𝒞 as a ‘homotopically universal’ construction. In this paper we will not need the full power of contractibility, but we will need the connectedness of this category. This says that any two choices of such (Re, Sing, η) can be connected by a zig-zag of Quillen homotopies.

7.6. Remark. The result in (c) is false if one does not assume that all the ℳ_{i}’s are additive. For an example, let R be the dga ℤ[e; de = 2]/(e^{4}) and let T be the dga ℤ/2[x; dx = 0]/(x^{2}), where e has degree 1 and x has degree 2. Let ℳ and 𝒩 be the categories of R- and T -modules, respectively. These turn out to be Quillen equivalent, but they cannot be linked by a zig-zag of Quillen equivalences between additive model categories. A verification of these claims can be found in [DS2, Section 8].

7.7. Additive presentations. We turn to the proof of Theorem 7.4(b). This will be deduced from the work of [D3] plus some purely formal considerations.

Let ℳ be a combinatorial model category. By [D3, Prop. 3.3], there is a small category 𝒞 and a functor 𝒞 → ℳ such that the induced map L: U𝒞 → ℳ is homotopically surjective (see [D3, Def. 3.1] for the definition). Then [D3, Prop. 3.2] shows that this fact implies there is a set of maps S in U𝒞 which the derived functor of L takes to weak equivalences, and such that the resulting map U𝒞/S → ℳ is a Quillen equivalence.

Now suppose that ℳ was also an additive model category. By examining the proof of [D3, Prop. 3.3] one sees that the 𝒞 constructed there is actually an additive category and the functor γ: 𝒞 → ℳ an additive functor taking values in the cofibrant objects (the category 𝒞 is a certain full subcategory of the cosimplicial objects over ℳ). By Theorem 7.4(a) there is an induced map F: U_{ad}𝒞 → ℳ. Again using [D3, Prop. 3.2], it will be enough to prove that this map is homotopically surjective.

Consider now the following sequence of adjoint pairs:

\xymatrix{Func(𝒞^{op}, s𝒮et) \ar[r]^{ℤ} & Func(𝒞^{op}, s𝒜b) \ar[r]^{Ad} \ar[l]^{U} & Func_{ad}(𝒞^{op}, s𝒜b) \ar[r]^{F} \ar[l]^{i} & ℳ \ar[l]^{Sing}}

The composite of the right adjoints is clearly the right adjoint of L, so the composite of the left adjoints is L. We have constructed things so that this composite is homotopically surjective, and we are trying to show that F is also homotopically surjective.

In the following lemma, note that the presheaf rX can be regarded as an object of either Func_{ad}(𝒞^{op}, 𝒜b) or Func(𝒞^{op}, 𝒮et). It will usually be clear from context which one we intend.

7.8. Lemma. If X ∈ 𝒞 then Ad(ℤ(rX)) ≅ rX. Or equivalently, Ad(ℤ(Ui(rX))) ≅ rX.

Proof. This follows from the fact that the two functors Func_{ad}(𝒞^{op}, 𝒜b) → 𝒜b given by F ↦ Func_{ad}(Ad(ℤ(rX)), F) and F ↦ Func_{ad}(rX, F) are both naturally isomorphic to F ↦ F (X).

Let G ∈ Func_{ad}(𝒞^{op}, s𝒜b). Let QG be the simplicial presheaf whose nth level is

∐_{rX_{n}→rX_{n−1}→···→rX_{0}→G_{n}}(rX_{n})

where the coproduct is in Func(𝒞^{op}, s𝒮et). The simplicial presheaf QG is treated in detail in [D2, Sec. 2.6], as it is a cofibrant-replacement functor for U𝒞. Likewise, let Q_{ad}G be the simplicial presheaf whose nth level is

⨁_{rX_{n}→rX_{n−1}→···→rX_{0}→G_{n}}(rX_{n})

where the coproduct is now in Func(𝒞^{op}, s𝒜b). The proof of [D3, Prop. 2.8] showing that Q is a cofibrant-replacement functor for U𝒞 adapts verbatim to show that Q_{ad} is a cofibrantreplacement functor for U_{ad}𝒞. Note that by Lemma 7.8 we have Q_{ad}G ≅ Ad(ℤ(Q(UiG))), since Ad and ℤ(−) are left adjoints and therefore preserve coproducts.

Finally we are in a position to conclude the

Proof of Theorem 7.4(b). We have reduced to showing that F: U_{ad}𝒞 → ℳ is homotopically surjective. Let Sing be the right adjoint of F. Then we must show that for every fibrant object X ∈ ℳ the induced map F Q_{ad}(Sing X) → X is a weak equivalence.

However, we have seen above that

F [Q_{ad} Sing X] ≅ F [Ad ℤ(QUi(Sing X))] ≅ L[QUi(Sing X)].

Recall that Ui Sing is the right adjoint to L. Since L: U𝒞 → ℳ is homotopically surjective we know LQ(Ui Sing X) → X is a weak equivalence in ℳ, so we are done.

8. Homotopy enrichments over Sp^{Σ}(s𝒜b)

In this section and the next we prove the main results stated in Section 1. Except for the work in the next section, the proofs are essentially the same as in [D4]—but they use Theorem 7.4 in place of [D4, Prop. 5.5].

8.1. Background on ring objects. If ℳ is a monoidal model category which is combinatorial and satisfies the monoid axiom, then by [SS1, Th. 4.1(3)] the category of monoids in ℳ has an induced model structure where the weak equivalences and fibrations are the same as those in ℳ. We’ll write ℛing[ℳ] for this model category. If 𝒩 is another such monoidal model category and L: ℳ ⇄ 𝒩: R is a Quillen pair which is weak monoidal in the sense of [SS3, Def. 3.6], then there is an induced Quillen map ℛing[ℳ] → ℛing[𝒩]. This is a Quillen equivalence if ℳ → 𝒩 was a Quillen equivalence and the units in ℳ and 𝒩 are cofibrant [SS3, Th. 3.12].

The adjunction 𝒮et_{∗} ⇄ 𝒜b is strong monoidal, and therefore induces strong monoidal Quillen functors Sp^{Σ}(s𝒮et_{∗}) ⇄ Sp^{Σ}(s𝒜b). It then follows that one gets a Quillen pair F: ℛing[Sp^{Σ}] ⇄ ℛing[Sp^{Σ}(s𝒜b)]: U. By the Eilenberg-Mac Lane ring spectrum associated to an R ∈ ℛing[Sp^{Σ}(s𝒜b)] we simply mean the ring spectrum UR.

8.2. Additive enrichments. Let ℳ be an additive, stable, combinatorial model category. By Theorem 7.4 there is a Quillen equivalence U_{ad}𝒞/S → ℳ for some small, additive category 𝒞 and some set of maps S in U_{ad}𝒞. The category U_{ad}𝒞/S is simplicial, left proper, and cellular, so using [Ho2, Sections 8, 9] we may form Sp^{Σ}(U_{ad}𝒞/S). Since U_{ad}𝒞/S is stable (since ℳ was), we obtain a zig-zag of Quillen equivalences

ℳ ⟵^{∼} U_{ad}𝒞/S ⟶^{∼} Sp^{Σ}(U_{ad}𝒞/S).

Applying ME_{0}(−, Sp^{Σ}(s𝒜b)) to this zig-zag gives a diagram of bijections by [D4, 3.14(d)].

The category U_{ad}𝒞 is a s𝒜b-model category, and therefore Sp^{Σ}(U_{ad}𝒞/S) is a Sp^{Σ}(s𝒜b)- model category by [Ho2, 8.3]. So Sp^{Σ}(U_{ad}𝒞/S) comes with a natural model enrichment by Sp^{Σ}(s𝒜b), as in [D4, Ex. 3.2]. We can transport this enrichment onto ℳ via the Quillen equivalences, and therefore get an element σ_{ℳ} ∈ ME_{0}(ℳ, Sp^{Σ}(s𝒜b)). Just as in [D4, Prop. 6.1], one shows (using Theorem 7.4) that this quasi-equivalence class does not depend on the choice of 𝒞, S, or the Quillen equivalence U_{ad}𝒞/S ⟶^{∼} ℳ.

We can now give the:

Proof of Theorem 1.3. We have just constructed the enrichment σ_{ℳ} . The proof that it is preserved by Quillen equivalences is exactly the same as in [D4, Prop. 6.2], but using Theorem 7.4.

Let X ∈ ℳ, and let X̃ be a cofibrant-fibrant object weakly equivalent to X. We write hEnd_{ad}(X) for any object in ℛing[Sp^{Σ}(s𝒜b)] having the homotopy type of σ_{ℳ} (X̃, X̃), and we’ll call this the additive homotopy endomorphism object of X. By [D4, Cors. 3.6, 3.7] this homotopy type depends only on the homotopy type of X and the quasi-equivalence class of σ_{ℳ} —and so it is a well-defined invariant of X and ℳ.

Proof of Proposition 1.4. This is entirely similar to the proof of [D4, Th. 1.4], but using Theorem 7.4(c).

Proof of Proposition 1.6. Same as the proof of [D4, Prop. 1.5].

Proof of Proposition 1.5. We know that there exists a zig-zag of Quillen equivalences ℳ ⟵^{∼} U_{ad}𝒞/S ⟶^{∼} Sp^{Σ}(U_{ad}𝒞/S). Therefore, using [D4, Thm. 1.4] and Proposition 1.4 we may as well assume ℳ = Sp^{Σ}(U_{ad}𝒞/S). This is an Sp^{Σ}(s𝒜b)-model category, and so for any object X we have a ring object \underline{ℳ}(X, X) in Sp^{Σ}(s𝒜b). The adjoint functors 𝒮et_{∗} ⇄ 𝒜b induce a strong monoidal adjunction F: Sp^{Σ}(s𝒮et_{∗}) ⇄ Sp^{Σ}(s𝒜b): U. The Sp^{Σ}(s𝒜b)-structure on ℳ therefore yields an induced Sp^{Σ}-structure as well (see [D4, Lem. A.5]). In this structure, the endomorphism ring spectrum of X is precisely U[\underline{ℳ}(X, X)]. Using [D4, Prop. 1.5], we know that this has the homotopy type of the ring spectrum hEnd(X), at least when X is cofibrant-fibrant. And Proposition 1.6 says that \underline{ℳ}(X, X) has the homotopy type of hEnd_{ad}(X). This is all we needed to check.

9. Chain enrichments

Proposition 1.6 says that if ℳ is a Sp^{Σ}(s𝒜b)-model category then one can compute hEnd_{ad}(X) using the Sp^{Σ}(s𝒜b)-structure. We would like to prove a similar result for Ch-model categories, where Ch denotes the model category of unbounded chain complexes of abelian groups. These are what arise most commonly in algebraic situations.

The monoidal model categories Sp^{Σ}(s𝒜b) and Ch can be connected by a zig-zag of weak monoidal Quillen equivalences, as described in [S]. This zig-zag can be used to translate enrichment-type information between these two categories. However, this is not as straightforward as one might expect; there are complications arising from the monoidal properties of the Dold-Kan equivalence between s𝒜b and ch_{+}, as analyzed in [SS3]. Our method for dealing with this requires some cumbersome machinery and gives a slightly weaker result than one would like. However, it is the best we can do at the moment.

9.1. Statement of the result. We give Ch the projective model structure, where weak equivalences are quasi-isomorphisms and fibrations are surjections. Recall again that a Ch-model category is a model category with compatible tensors, cotensors, and enrichments over Ch satisfying an analogue of SM7; see Section 2. For X, Y in ℳ, we denote the enriched hom-object in Ch by \underline{ℳ}_{Ch}(X, Y).

Note that a Ch-model category is automatically additive and stable. See Corollary 6.9 for the additivity, and [SS2, 3.5.2] or [GS, 3.2] for stability.

Recall from [S] that there are two Quillen equivalences

\xymatrix{Sp^{Σ}(ch_{+}) \ar[r]^{D} & Ch \ar[l]^{R} & and & Sp^{Σ}(ch_{+}) \ar[r]^{L} & Sp^{Σ}(s𝒜b) \ar[l]^{ν}}

in which (D, R) is strong monoidal and (L, ν) is weak monoidal. These induce Quillen equivalences between the corresponding model categories of rings:

ℛing(Sp^{Σ}(ch_{+})) ⇄^{∼} 𝒟G𝒜 ℛing(Sp^{Σ}(ch_{+})) ⇄^{∼} ℛing(Sp^{Σ}(s𝒜b)).  (9.2)

In the first equivalence of (9.2) the left and right adjoints are just the restrictions of D and R, as these were strong monoidal. In the second, the right adjoint is just ν again, but the left adjoint is more complicated; see [SS3, 3.3].

Let \underline{ν} and \underline{D} denote the derived functors of ν and D from (9.2), and write Θ′ = \underline{D} \underline{ν}. So Θ′ is a functor

ℋo(ℛing(Sp^{Σ}(s𝒜b))) → ℋo(𝒟G𝒜).

Let ℳ be a stable, combinatorial, additive model category and let X ∈ ℳ. We have shown how to associate to X an object hEnd_{ad}(X) ∈ ℛing(Sp^{Σ}(sAb)). By applying Θ′ we get the homotopy endomorphism dga of X. Denote this as hEnd_{dga}(X) = Θ′[hEnd_{ad}(X)].

The goal for this section is to prove Proposition 1.7. We restate the result here for the convenience of the reader.

9.3. Proposition. Suppose that ℳ is a combinatorial Ch-model category, and that ℳ has a generating set of compact objects. Let X ∈ ℳ be cofibrant and fibrant. Then the dga \underline{ℳ}_{Ch}(X, X) is quasi-isomorphic to hEnd_{dga}(X)

Proposition 9.3 will be proven by reducing from a Ch-model category to a Sp^{Σ}(s𝒜b)- model category and then applying results of Section 8. The reduction from Ch to Sp^{Σ}(ch_{+}) will be simple because of the strong monoidal equivalence between these two categories. The following proposition provides the reduction from Sp^{Σ}(ch_{+}) to Sp^{Σ}(s𝒜b). This is where all the enriched category theory from Sections 2 through 5 is needed. Recall that for a general 𝒟-model category 𝒩 we denote the morphism object in 𝒟 by \underline{𝒩}_{𝒟} (X, Y).

9.4. Proposition. Let ℳ be a combinatorial Sp^{Σ}(ch_{+})-model category with a generating set of compact objects. Let X ∈ ℳ be a cofibrant-fibrant object. Then there exists

(i) a combinatorial, Sp^{Σ}(s𝒜b)-model category 𝒩,

(ii) a zig-zag of Quillen equivalences between ℳ and 𝒩, where the intermediate model categories are all additive, and

(iii) a cofibrant-fibrant object Y ∈ 𝒩

such that Y is taken to X by the derived functors of the Quillen equivalences and \underline{ν}[\underline{𝒩}_{Sp^{Σ}(s𝒜b )}(Y, Y)] is weakly equivalent to \underline{ℳ}_{Sp^{Σ}(ch_{+})}(X, X).

Proof. This is a special case of Proposition 5.2 and Corollary 5.3. We need to verify the properties for 𝒞 = Sp^{Σ}(ch_{+}) and 𝒟 = Sp^{Σ}(s𝒜b) stated just prior to Proposition 5.2, with (F, G) replaced by (L, ν). Axioms (QI1-2) for 𝒞 follow from [S, 3.2, 3.3]. The fact that (L, ν) is a weak monoidal Quillen equivalence is given in [S, 4.3]. All the other conditions are easy exercises, but see also Example 4.13 for more information.

Using the above proposition, we can complete the following:

Proof of Proposition 9.3. Let ℳ be a combinatorial Ch-model category with a generating set of compact objects. Let 𝒞 = Sp^{Σ}(ch_{+}). Using the strong monoidal adjunction (D, R), ℳ becomes a 𝒞-model category via the definitions Z ⊗ c = Z ⊗ D(c), Z^{c} = Z^{Dc}, and \underline{ℳ}_{𝒞}(W, Z) = R[\underline{ℳ}_{Ch}(W, Z)] where W, Z ∈ ℳ and c ∈ 𝒞. See [D4, Lem. A.5].

Now we apply Proposition 9.4 to ℳ with this 𝒞-model structure to construct 𝒩 and Y. By Proposition 1.4, the additive homotopy endomorphism spectra corresponding to X and Y are weakly equivalent. Let 𝒟 = Sp^{Σ}(s𝒜b). Since 𝒩 is a 𝒟-model category, we have by Proposition 1.6 that hEnd_{ad}(Y) is weakly equivalent to \underline{𝒩}_{𝒟} (Y, Y). So we have

hEnd_{dga}(X) = Θ′[hEnd_{ad}(X)] ≃ Θ′[hEnd_{ad}(Y)] ≃ \underline{D} \underline{ν}[\underline{𝒩}_{𝒟} (Y, Y)]

(recalling that Θ′ = \underline{D} \underline{ν}).

But 𝒩 and Y were chosen in such a way that we have \underline{ν}[\underline{𝒩}_{𝒟} (Y, Y)] ≃ \underline{ℳ}_{𝒞}(X, X). So in fact

hEnd_{dga}(X) ≃ \underline{D} [\underline{ℳ}_{𝒞}(X, X)] = \underline{D} \underline{R}[\underline{ℳ}_{Ch}(X, X)] ≃ \underline{ℳ}_{Ch}(X, X).

A. Homotopy theory of 𝒞I-categories

The present section reviews and expands on results from [SS3]. In particular, [SS3] often states results in settings which are extremely general and therefore require somewhat awkward hypotheses. Here we will specialize, replacing those hypotheses with conditions more readily checked in practice.

We assume that 𝒞 is a combinatorial, symmetric monoidal model category. Also, 𝒞 is assumed to satisfy the monoid axiom of [SS1, 3.3]. We’ll refer to those conditions as our ‘standing assumptions’. Finally, we will sometimes require the following two conditions as well:

(QI1) For any cofibrant object A ∈ 𝒞 and any weak equivalence X → Y, the map A⊗X → A ⊗ Y is also a weak equivalence.

(QI2) Suppose A ↣ B is a cofibration, and X is any object. Then for any map A⊗X → Z, the map from the homotopy pushout of B ⊗ X ⟵ A ⊗ X → Z to the pushout is a weak equivalence.

The abbreviation (QI) is for ‘Quillen invariance’, as these conditions will be used to check what [SS3, 3.11] calls Quillen invariance for modules.

A.1. Example. The category ch_{+} of non-negatively graded chain complexes with tensor product and its usual ‘projective’ model structure satisfies (QI1-2). It follows from [S, 3.2, 3.3] that Sp^{Σ}(ch_{+}) also satisfies (QI1-2). Typically, these axioms will follow from the existence of an ‘injective’ model structure for ℳ in which all objects are cofibrant, provided such a model structure is a Quillen module over the corresponding projective version.

Let I be a set. We assume the reader is familiar with the notion of 𝒞I-category (a category enriched over 𝒞 with object set I) from [Bo, 6.2]. If 𝒪 is a 𝒞I-category, then the category of right 𝒪-modules (contravariant 𝒞-functors from 𝒪 to 𝒞) is defined in [Bo, 6.2]; see also [SS3, Section 6].

A.2. Proposition. Let 𝒪 be a 𝒞I-category.

(a) The category Mod- 𝒪 has a model category structure in which the weak equivalences and fibrations are defined objectwise.

(b) Let 𝒪 → ℛ be a map of 𝒞I-categories. Then there is a Quillen map Mod- 𝒪 → Mod- ℛ in which the right adjoint is restriction.

(c) If 𝒪 → ℛ is a weak equivalence and 𝒞 satisfies (QI1-2), then the above Quillen map is a Quillen equivalence.

Proof. Part (a) is [SS3, 6.1(1)]. For (b) we need only construct the left adjoint, as the restriction clearly preserves fibrations and trivial fibrations. This construction is given in the paragraph above [SS3, 6.1]. Denote this left adjoint by X ↦ X ⊗_{𝒪} ℛ.

Part (c) requires a little work. First, for any i ∈ I and A ∈ 𝒞 let A ⊗ F r_{i}(𝒪) be the ‘free 𝒪-module generated by A at spot i’. This is defined by j ↦ A ⊗ 𝒪(j, i). It is easy to see that we have the adjunction Mod- 𝒪(A ⊗ F r_{i}(𝒪), X) ≅ 𝒞(A, X(i)). From this it immediately follows that

[A ⊗ F r_{i}(𝒪)] ⊗_{𝒪} ℛ ≅ A ⊗ F r_{i}(ℛ).

As a another consequence of the adjunction, observe that Mod- 𝒪 is cofibrantly-generated and the generating cofibrations are maps of the form A ⊗ F r_{i}(𝒪) → B ⊗ F r_{i}(𝒪) where A ↣ B is a generating cofibration of 𝒞.

By [SS3, 6.1(2)], to prove (c) it suffices to check that for any cofibrant 𝒪-module N the natural map N → U[N ⊗_{𝒪} ℛ] is a weak equivalence, where U is the restriction Mod- ℛ → Mod- 𝒪. Let G denote the composite functor X ↦ U[X ⊗_{𝒪} ℛ], so that we are concerned with the natural transformation Id → G. Note that when X = A ⊗ F r_{i}(𝒪) we have G(X) = A ⊗ F r_{i}(ℛ). If A is cofibrant, the map X → G(X) is an objectwise weak equivalence because 𝒪 → ℛ is (this uses (QI1)).

Apply the small object argument to factor ∅ → N as a cofibration followed by a trivial fibration. This gives us a (possibly transfinite) sequence of cofibrations

∅ = W_{0} ↣ W_{1} ↣ W_{2} ↣ ⋯

in which W_{i+1} is obtained from W_{i} by a pushout diagram

\xymatrix{∐_{j} A_{j} ⊗ F r_{j}(𝒪) \ar[r] \ar[d] & W_{i} \ar[d] \\ ∐_{j} B_{j} ⊗ F r_{j}(𝒪) \ar[r] & W_{i+1},}

together with a trivial fibration W_{∞} = colim_{i} W_{i} → N. Since N is cofibrant, N is a retract of W_{∞}. So it will suffice to show that W_{∞} → G(W_{∞}) is a weak equivalence, as N → GN is a retract of this map.

We first prove that if W_{i−1} → G(W_{i−1}) is a weak equivalence then the same is true of W_{i} → G(W_{i}). To see this, note that we have the following diagram:

[figure: ∐ B_{j} ⊗ F r_{j}(𝒪); ∼; ∐ A_{j} ⊗ F r_{j}(𝒪); ∼; W_{i−1}; ∼; ∐ G(B_{j} ⊗ F r_{j}(𝒪)); ∐ G(A_{j} ⊗ F r_{j}(𝒪)); G(W_{i−1}).]

The pushout of the top row is W_{i}, and of the bottom row is G(W_{i}) (the latter follows because G preserves colimits). Note that G(A_{j} ⊗ F r_{j}(𝒪)) → G(B_{j} ⊗ F r_{j}(𝒪)) is a cofibration, as it is just the map A_{j} ⊗ F r_{j}(ℛ) → B_{j} ⊗ F r_{j}(ℛ). It follows that G(W_{i−1}) → G(W_{i}) is a cofibration.

Certainly the above diagram induces a weak equivalence of homotopy pushouts. We claim these homotopy pushouts are weakly equivalent to the corresponding pushouts. This is an objectwise question, since pushouts, homotopy pushouts, and weak equivalences in the module category are all determined objectwise. The claim for the top row then follows directly from (QI2). The claim for the bottom row is similar, but uses the identification G(B_{j} ⊗ F r_{j}(𝒪)) = B_{j} ⊗ F r_{j}(ℛ), etc.

Thus, we have shown that W_{i} → G(W_{i}) is a weak equivalence whenever W_{i−1} → G(W_{i−1}) is so. It is trivial that W_{0} → G(W_{0}) is a weak equivalence. The result now follows by a transfinite induction, using [Hi, 17.9.1] to pass the weak equivalences to the limit ordinals. One again uses that G preserves colimits.

A.3. Proposition. Let I be a fixed set, and let L: 𝒞 ⇄ 𝒟: R be a weak monoidal Quillen pair (see Section 2.4) where both 𝒞 and 𝒟 satisfy our standing assumptions.

(a) The category 𝒞I − 𝒞at (and likewise 𝒟I − 𝒞at) has a model category structure in which weak equivalences and fibrations are defined objectwise.

(b) There is a Quillen map 𝒞I − 𝒞at → 𝒟I − 𝒞at in which the right adjoint is ‘apply R objectwise’. The left adjoint will be denoted L^{𝒟I} .

(c) Suppose 1_{𝒞} and 1_{𝒟} are cofibrant. If 𝒪 is a cofibrant 𝒞I-category then there are weak equivalences L[𝒪(i, j)] → (L^{𝒟I} 𝒪)(i, j) for every i, j ∈ I. These are adjoint to the maps provided by the adjunction unit 𝒪 → R(L^{𝒟I} 𝒪).

(d) Suppose (L, R) is a Quillen equivalence and 1_{𝒞}, 1_{𝒟} are cofibrant. Then the induced Quillen map 𝒞I − 𝒞at → 𝒟I − 𝒞at is also a Quillen equivalence.

Proof. Part (a) is [SS3, 6.3(1)]. For part (b) we argue as follows. Recall the category 𝒞I − 𝒢raph from [SS3, 6.1], and that this category comes equipped with a monoidal product ⊗. A 𝒞I-category is precisely a monoid with respect to this tensor product. The existence of the desired left adjoint follows from Lemma A.4 below. As the right adjoint obviously preserves fibrations and trivial fibrations, we have a Quillen pair.

For part (c), note that there is a Quillen map 𝒞I −𝒢raph → 𝒞I −𝒞at in which the right adjoint is the forgetful functor. The model structure on 𝒞I − 𝒞at is ‘created’ by these adjoint functors from the cofibrantly-generated model structure on 𝒞I − 𝒢raph. [SS3, 6.4(1)] proves the desired claim in the case 𝒪 is a cell complex, but since any cofibrant object is a retract of a cell complex one immediately obtains the more general statement.

Finally, we prove (d). Note that since the functor 𝒟I − 𝒞at → 𝒞I − 𝒞at is just ‘apply R objectwise’, a map of fibrant objects X → Y in 𝒟I − 𝒞at is a weak equivalence if and only if RX → RY is a weak equivalence in 𝒞I − 𝒞at. So by Lemma A.5 below, we only need to show that if 𝒪 is a cofibrant 𝒞I-category and L^{𝒟I} 𝒪 ⟶^{∼} 𝒜 is a fibrant replacement in 𝒟I − 𝒞at, then 𝒪 → R𝒜 is a weak equivalence.

Since weak equivalences are detected objectwise, we must check that 𝒪(i, j) → R[𝒜(i, j)] is a weak equivalence for every i, j ∈ I. But 𝒪 is cofibrant, so each 𝒪(i, j) is cofibrant in 𝒞 (see [SS3, 6.3(2)] —this uses that 1_{𝒞} is cofibrant). And since 𝒜 is fibrant, each 𝒜(i, j) is fibrant. Using the Quillen equivalence (L, R), we are therefore reduced to checking that L[𝒪(i, j)] → 𝒜(i, j) is a weak equivalence. But we are really looking at the composite

L[𝒪(i, j)] → [L^{𝒟I} 𝒪](i, j) → 𝒜(i, j).

The second map was assumed to be a weak equivalence, and the first map is a weak equivalence by part (c). So we are done.

In this proof we used the following two lemmas.

A.4. Lemma. Let 𝒞 be a monoidal category which is complete and co-complete. Assume that for any X ∈ 𝒞 the functors X ⊗ (−) and (−) ⊗ X preserve filtered colimits and countable coproducts. Then

(a) The category of monoids in 𝒞 is co-complete.

(b) If ℬ is another monoidal category, and L: ℬ ⇄ 𝒞: R is an adjunction where R is weak monoidal, then R induces a functor 𝒞−ℳonoid → ℬ−ℳonoid and this functor has a left adjoint.

Proof. Let T: 𝒞 → 𝒞 be the ‘free algebra’ monad, where

T (X) = 1 ⨿ X ⨿ (X ⊗ X) ⨿ · · · .

(The fact that this is a monad uses the assumption that tensoring preserves countable coproducts). The monoids in 𝒞 are precisely the T -algebras. Our assumptions imply that T preserves filtered colimits, so [Bo, 4.3.6] implies that 𝒞 − ℳonoid is co-complete.

Part (b) is an immediate consequence of (a) and [Bo, 4.5.6].

A.5. Lemma. [Ho1, 1.3.16] Let L: ℳ ⇄ 𝒩: R be a Quillen pair. Assume the following two conditions hold:

(i) If X and Y are fibrant objects in 𝒩, a map X → Y is a weak equivalence whenever RX → RY is a weak equivalence.

(ii) For every cofibrant object A ∈ ℳ and every fibrant replacement LA ⟶^{∼} Z in 𝒩, the composite map A → RLA → RZ is a weak equivalence.

Then (L, R) is a Quillen equivalence.

Here is the final result we will need:

A.6. Proposition. Again assume that L: 𝒞 ⇄ 𝒟: R is a weak monoidal Quillen pair, where 𝒞 and 𝒟 satisfy our standing assumptions. Also assume that 1_{𝒞} and 1_{𝒟} are cofibrant.

(a) If 𝒜 is a 𝒟I-category then there is a Quillen map Mod- R𝒜 → Mod- 𝒜 in which the right adjoint is ‘apply R objectwise’.

(b) Let 𝒪 be a 𝒞I-category. Then there is a Quillen map Mod- 𝒪 → Mod-(L^{𝒟I} 𝒪) in which the right adjoint is the composition of ‘applying R objectwise, then restricting across 𝒪 → R(L^{𝒟I} 𝒪)’.

(c) If L: 𝒞 ⇄ 𝒟: R is a Quillen equivalence and 𝒪 is a cofibrant 𝒞I-category, then Mod- 𝒪 → Mod-(L^{𝒟I} 𝒪) is also a Quillen equivalence.

Proof. If X is an object in the functor category 𝒟^{I}, let T_{𝒜} X ∈ 𝒟^{I} be the functor j ↦ ∐_{j} X(j) ⊗ 𝒜(−, j). Note that this is a monad in an obvious way, and that the T_{𝒜} -algebras are precisely the 𝒜-modules. We have the diagram of categories

\xymatrix{𝒞^{I} & 𝒟^{I} \ar[l]^{R} \\ Mod-(R𝒜) \ar[u] & Mod- 𝒜 \ar[l]^{R} \ar[u]}

where the vertical maps are forgetful functors. By [Bo, 4.5.6] the map Mod- 𝒜 → Mod-(R𝒜) has a left adjoint, since Mod- 𝒜 is cocomplete. This clearly gives a Quillen pair.

For (b) we use the composite of the two Quillen maps

Mod- 𝒪 → Mod-(RL^{𝒟I} 𝒪) → Mod-(L^{𝒟I} 𝒪).

The first is provided by Proposition A.2(b), induced by the map 𝒪 → RL^{𝒟I} 𝒪. The second comes from (a) of the present result.

Finally, part (c) is just [SS3, 6.5(1)].

References

[AR] J. Adamek and J. Rosicky, Locally presentable and accessible categories, London Math. Society Lecture Note Series 189, Cambridge University Press, 1994.

[Bo] F. Borceux, Handbook of categorical algebra 2: Categories and structures, Cambridge University Press, 1994.

[BF] A. K. Bousfield and E. M. Friedlander, Homotopy theory of Γ-spaces, spectra, and bisimplicial sets, Geometric Applications of Homotopy Theory II, Lect. Notes. in Math., vol. 658, Springer-Verlag, New York, 1978, pp. 80–130.

[Dr] V. Drinfeld, Quotients of dg-categories, J. Algebra 272 (2004), no. 2, 643–691.

[D1] D. Dugger, Replacing model categories with simplicial ones, Trans. Amer. Math. Soc. 353, no. 12 (2001), 5003–5027.

[D2] D. Dugger, Universal homotopy theories, Adv. Math. 164, no. 1 (2001), 144–176.

[D3] D. Dugger, Combinatorial model categories have presentations, Adv. Math. 164, no. 1 (2001), 177-201.

[D4] D. Dugger, Spectral enrichments of model categories, Homology Homotopy Appl. 8 (1), 2006.

[DS1] D. Dugger and B. Shipley, K-theory and derived equivalences, Duke Math. J. 124 (2004), no. 3, 587–617.

[DS2] D. Dugger and B. Shipley, Topological equivalences for differential graded algebras, Adv. Math. 212 (2007), no. 1, 37–61.

[DS3] D. Dugger and B. Shipley, A curious example of two model categories and some associated differential graded algebras, in preparation.

[DK] W.G. Dwyer and D.M. Kan, Function complexes in homotopical algebra, Topology 19 (1980), 427–440.

[GS] J. P. C. Greenlees and B. Shipley, An algebraic model for rational torus-

        equivariant spectra. Preprint, 2004. http://www.math.uic.edu/~bshipley/

[Hi] P. Hirschhorn, Model Categories and Their Localizations, Mathematical Surveys and Monographs, vol. 99, Amer. Math. Soc., 2003.

[Ho1] M. Hovey, Model categories, Mathematical Surveys and Monographs, 63, American Mathematical Society, Providence, RI, 1999, xii+209 pp.

[Ho2] M. Hovey, Spectra and symmetric spectra in general model categories, J. Pure Appl. Algebra 165 (2001), no. 1, 63–127.

[HSS] M. Hovey, B. Shipley, and J. Smith, Symmetric spectra, J. Amer. Math. Soc. 13 (2000) 149–208.

[ML] S. Mac Lane, Categories for the working mathematician, Graduate Texts in Math. 5, Springer, New York-Berlin, 1971.

[SS1] S. Schwede and B. Shipley, Algebras and modules in monoidal model categories, Proc. London Math. Soc. 80 (2000) 491-511.

[SS2] S. Schwede and B. Shipley, Stable model categories are categories of modules, Topology 42 (2003), 103-153.

[SS3] S. Schwede and B. Shipley, Equivalences of monoidal model categories, Algebr. Geom. Topol. 3 (2003), 287–334.

[S] B. Shipley, Hℤ-algebra spectra are differential graded algebras, Amer. Jour. Math. 129 (2007), 351–379.

[T] B. Toën, The homotopy theory of dg-categories and derived Morita theory, Invent. Math. 167 (2007), 615–667.

Department of Mathematics University of Oregon Eugene, OR 97403

Department of Mathematics University of Illinois at Chicago Chicago, IL 60607

Email: ddugger@math.uoregon.edu
      bshipley@math.uic.edu

This article may be accessed at http://www.tac.mta.ca/tac/ or by anonymous ftp at

ftp://ftp.tac.mta.ca/pub/tac/html/volumes/18/15/18-15.{dvi,ps,pdf}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX2e strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: mbarr@barrs.org

Transmitting editors.

Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown (at) btinternet.com Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it Valeria de Paiva, Xerox Palo Alto Research Center: paiva@parc.xerox.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk G. Max Kelly, University of Sydney: maxk@maths.usyd.edu.au Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, University of Western Sydney: s.lack@uws.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Brooke Shipley, University of Illinois at Chicago: bshipley@math.uic.edu James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca